### v2

# Asymptotically Distribution-free Goodness-of-Fit Testing for Point Processes

Abstract.
Consider an observation of a multivariate temporal point process $N$ with law $\mathcal{P}$ on the time interval $[0,T]$. To test the null hypothesis that $\mathcal{P}$ belongs to a given parametric family, we construct a convergent compensated counting process to which we apply an innovation martingale transformation. We prove that the resulting process converges weakly to a standard Wiener process. Consequently, taking a suitable functional of this process yields an asymptotically distribution-free goodness-of-fit test for point processes. For several standard tests based on the increments of this transformed process, we establish consistency under alternative hypotheses. Finally, we assess the performance of the proposed testing procedure through a Monte Carlo simulation study and illustrate its practical utility with two real-data examples.
Keywords. Point process $\circ$ goodness-of-fit testing $\circ$ innovation martingale transformation $\circ$ Hawkes process $\circ$ Monte Carlo simulation
Affiliations. JB, UC and RL are with the Dept. of Quantitative Economics, University of Amsterdam, Roetersstraat 11, 1018 WB Amsterdam, the Netherlands. RL is also with Eurandom, Eindhoven University of Technology, Den Dolech 2, 5612 AE Eindhoven, the Netherlands; and with CentER, Tilburg University,Warandelaan 2, 5037 AB Tilburg, the Netherlands. The research of JB and RL is funded in part by the Netherlands Organization for Scientific Research under an NWO VICI grant (2020–2027).
Email. J.R.Baars@uva.nl, S.U.Can@uva.nl, R.J.A.Laeven@uva.nl.
Date: June 25, 2025.
1. Introduction
Consider a $d$-variate temporal point process $N=(N^{(1)},\ldots,N^{(d)})$ depending on time $t\in\mathbb{R}$. Suppose that $N$ is modeled through a parametric null hypothesis $\mathscr{F}_{\Theta}:=\{N_{\theta}:\theta\in\Theta\}$, where $\Theta\subset\mathbb{R}^{m}$ is a finite-dimensional parameter space. Under the null hypothesis $H_{0}$, it holds that $N\stackrel{{\scriptstyle d}}{{=}}N_{\theta_{0}}$ for some $\theta_{0}\in\Theta$. In practice, $\theta_{0}$ is typically unknown, and has to be estimated using an observation $\omega_{T}=\{t_{1},\ldots,t_{N(T)}\}$ of the point process on the time interval $[0,T]$, leading to an estimated parameter vector $\hat{\theta}_{T}$. If maximum likelihood estimation is employed, the statistical properties of $\hat{\theta}_{T}$ are well-established [42, 48, 49].
The present work is concerned with goodness-of-fit testing for temporal point processes. Given an observation $\omega_{T}$ of the point process on the time interval $[0,T]$, we are interested in testing the hypothesis $H_{0}:N\stackrel{{\scriptstyle d}}{{=}}N_{\theta}$ for some $\theta\in\Theta$ against the alternative $H_{1}:N\stackrel{{\scriptstyle d}}{{\neq}}N_{\theta}$ for all $\theta\in\Theta$. In goodness-of-fit problems, the test statistic $S$ often has a distribution that is challenging to calculate and that may depend on the null hypothesis or on the true parameter, fragmenting the theory. As a result, one has to resort to ad hoc bootstrap methods to estimate critical values for each specific application. To address this issue, a test statistic, or test process, is needed that, under the null hypothesis, has a distribution that is independent of both $\mathscr{F}_{\Theta}$ and the true parameter $\theta_{0}$.
For temporal point processes, a common approach is to apply a time transformation $\left(\Lambda(t_{i})\right)_{i\in[N(T)]}$, where $[n]:=\{1,\ldots,n\}$ and $\Lambda(t)=\int_{0}^{t}\lambda(s)\ \mathrm{d}s$ is the compensator of the process, i.e., the time-integrated conditional intensity. According to the random time change theorem [16], Theorem 7.4.I, under the law $\mathcal{P}_{\theta_{0}}$ of $N_{\theta_{0}}$ the transformed arrival times $\Lambda_{\theta_{0}}(t_{i})$ follow a Poisson process of unit rate. Since $\theta_{0}$ is unknown, typically the goodness-of-fit of a candidate model is assessed by testing whether the transformed interarrival times $\Lambda_{\hat{\theta}_{T}}(t_{i})-\Lambda_{\hat{\theta}_{T}}(t_{i-1})$ are standard exponential. In such tests, the estimation uncertainty $\hat{\theta}_{T}\neq\theta_{0}$, which causes a discrepancy between $\Lambda_{\hat{\theta}_{T}}$ and $\Lambda_{\theta_{0}}$, is generally ignored; see [3, 11, 13, 22, 34, 35, 39, 57, 58].
In recent decades, several goodness-of-fit testing procedures for temporal point processes have been proposed, such as random thinning [43], random superposition [8], super-thinning [14], and random-time-change-based methods [61, 53]. However, these methods are based on results that hold only under the true distribution $\mathcal{P}_{\theta_{0}}$. In practice, since $\theta_{0}$ is unknown and typically replaced by the estimator $\hat{\theta}_{T}$, these tests fail to account for the uncertainty introduced by estimation.
Ignoring estimation uncertainty has important consequences: the distribution of the test statistic under the null hypothesis still depends on the specific model class $\mathscr{F}_{\Theta}$ and the true parameter $\theta_{0}$, even in the limit as $T\to\infty$. There is no theoretical justification for ignoring the estimation uncertainty $\hat{\theta}_{T}\neq\theta_{0}$. In fact, even if the data is generated under the law $\mathcal{P}_{\theta_{0}}$, the estimator $\hat{\theta}_{T}$ will, by construction, fit the observed data better than the true parameter $\theta_{0}$. Consequently, for a test statistic $S(\omega_{T},\theta)$ depending on both the observation $\omega_{T}$ and the parameter $\theta$, it is natural to expect that under $\mathcal{P}_{\theta_{0}}$, the value of $S(\omega_{T},\hat{\theta}_{T})$ will appear “less extreme” than $S(\omega_{T},\theta_{0})$. Ignoring estimation uncertainty leads to undersized tests with reduced power. In applied work, it is common to fit a model to data and then attempt to demonstrate that the model fits well by using a goodness-of-fit test that fails to reject the null hypothesis. Ignoring estimation uncertainty in this context leads to overly conservative tests (i.e., with too low power), potentially giving a false sense of confidence in an otherwise imperfect model.
In recent years, efforts have been made to address this issue. Asymptotically correct goodness-of-fit tests have been developed for the specific cases of self-exciting temporal point processes [19] and exponential Hawkes processes [7], although some of these results lack theoretical guarantees. Additionally, sparsity tests for the connectivity matrices of multivariate Hawkes processes have been introduced [38]. Further progress has been made by [51], who mitigate the bias caused by using the plugged-in estimated parameter in scenarios with a large number of sample paths. However, their results do not directly apply to the case where only a single sample path is observed. Moreover, asymptotically correct goodness-of-fit tests have been proposed for cases where the temporal dynamics are partially known a priori [15, 52]. Nevertheless, none of these recent works presents an asymptotically exact goodness-of-fit testing procedure for a single observation of a general point process. To the best of our knowledge, no existing approach directly addresses this challenge, leaving a significant gap in the field.
In this paper, we address the limitations of existing goodness-of-fit procedures by developing an asymptotically distribution-free testing method for a broad class of point processes. Specifically, based on the observation $\omega_{T}$ we construct a test process on a bounded interval $[0,\tau]$, which converges weakly to a standard Wiener process, as $T\to\infty$. Consequently, the limiting process has an asymptotic distribution that is independent of both the null hypothesis and the true parameter $\theta_{0}$. Functionals of this test process can be employed as a test statistic. Since the weak limit of the test process is a standard Wiener process that does not depend on the model class $\mathscr{F}_{\Theta}$ or the true parameter $\theta_{0}$, the asymptotic distributions of these functionals are also independent of $\mathscr{F}_{\Theta}$ and $\theta_{0}$. This means that their critical values only need to be tabulated once, making them applicable across a wide range of testing problems — an important practical advantage.
To derive the asymptotically distribution-free test process, we first establish a functional central limit theorem (FCLT) for the counting process $N_{\theta_{0}}$ compensated by $\Lambda_{\hat{\theta}_{T}}$. The resulting limit is the sum of a Wiener process and a linear drift with a random coefficient, which still depends on the true parameter $\theta_{0}$ and the model class $\mathscr{F}_{\Theta}$. However, the specific structure of this limit enables the application of an innovation martingale transformation $\mathscr{T}_{\theta_{0}}$, originally introduced by Khmaladze in [27, 28, 29]. After transformation, we obtain a process that, as we formally establish, converges to a standard Wiener process, as $T\to\infty$.
This paper contributes to the literature by introducing a new framework for goodness-of-fit testing for point processes with unknown finite-dimensional parameters. The resulting test process is asymptotically distribution-free, enabling inference without knowledge of nuisance parameters. Our main theoretical contribution is the FCLT for the compensated empirical counting process, obtained by combining martingale arguments with a careful analysis of the difference between the true and empirical compensators, and by application of an innovation martingale transformation. In developing this framework, we also establish a Cramér–Wold device for stochastic processes and derive an FCLT for multivariate martingales — results that may be of independent interest. The proposed tests are based on classical functionals of the limiting process, remain consistent under general alternatives, and account for estimation uncertainty. Compared to existing methods, this leads to improved reliability and greater power. We demonstrate these advantages through simulation studies and applications to real data.
The remainder of this paper is organized as follows. Section 2 introduces the setup and assumptions. Section 3 derives the functional central limit theorem for the compensated empirical counting process after applying an observable innovation martingale transformation. Section 4 focuses on goodness-of-fit testing: Section 4.1 proposes asymptotically distribution-free test statistics, while Section 4.2 discusses biases in random-time-change-based methods. Section 5 establishes consistency under general alternatives. Section 6 presents simulation results comparing our method to existing approaches. Section 7 illustrates the method’s performance in real-data applications. Section 8 outlines possible extensions.
Appendix
Several technical assumptions are collected in the appendix.
2. Model description and assumptions
Let $N=(N^{(1)},\ldots,N^{(d)})$ be a $d$-variate point process on $\mathbb{R}$, i.e., a random discrete $[d]$-marked subset of $\mathbb{R}$. Denote the law of $N$ by $\mathcal{P}$. A realization of $N$ consists of event times $\omega:=\{\ldots,t_{-2},t_{-1},t_{0},t_{1},t_{2},\ldots\}\subset\mathbb{R}$ along with coordinates $c_{i}\in[d]$; here, $t_{i}<t_{i+1}$ and $t_{0}<0\leqslant t_{1}$. The realization on $[0,T]$ is then denoted by $\omega_{T}=\omega\cap[0,T]=\{t_{1},\ldots,t_{N(T)}\}$, where $N(T)$ denotes the number of events in the interval $[0,T]$. We assume that $\omega$ has no limit point.
We are mostly interested in the situation in which $N\stackrel{{\scriptstyle d}}{{=}}N_{\theta_{0}}$, where $N_{\theta_{0}}$ belongs to the parametric family $\mathscr{F}_{\Theta}:=\{N_{\theta}:\theta\in\Theta\}$, with $\Theta\subset\mathbb{R}^{m}$. Here, =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution. Denote the law of $N_{\theta}$ by $\mathcal{P}_{\theta}$. We use this framework to test the parametric null hypothesis
$H_{0}:N\stackrel{{\scriptstyle d}}{{=}}N_{\theta}\text{ for some }\theta\in\Theta,$ (1)
using an observation $\omega_{T}$ of $N$ on the time interval $[0,T]$.
We start with assumptions on $N_{\theta}$ and on $\Theta$.
Assumptions A.
  1. (i)
    $\Theta\subset\mathbb{R}^{m}$ is open, convex and bounded.
  2. (ii)
    For each $\theta\in\Theta$, $N_{\theta}$ is a stationary and ergodic point process, absolutely continuous w.r.t. the unit Poisson process on $[0,T]$, for each $T\in(0,\infty)$.
  3. (iii)
    For each $\theta\in\Theta$, $N_{\theta}$ is an orderly point process, i.e., $\mathbb{P}(N_{\theta}[0,\delta]\geqslant 2)=o(\delta)$, as $\delta\downarrow 0$.
By Assumptions A, it follows that, for each $\theta\in\Theta$, the probabilistic behavior of $N_{\theta}$ can be characterized by a conditional intensity function $\lambda_{\theta}:\mathbb{R}_{+}\to\mathbb{R}_{+}^{d}:t\mapsto\lambda_{\theta}(t)$; see [16], Proposition 7.3.IV. Denote its components by $\lambda_{\theta}^{(k)}$, $k\in[d]$. Let
$\mathcal{H}_{t}:=\sigma(N_{\theta}(s):s\in(-\infty,t]).$
Then $\lambda_{\theta}$ can be taken to be any $(\mathcal{H}_{t})_{t\in\mathbb{R}}$-predictable function such that
$\lambda_{\theta}(t)=\lim_{\Delta t\downarrow 0}\frac{1}{\Delta t}\mathbb{P}\left(N_{\theta}[t,t+\Delta t)>0|\mathcal{H}_{t}\right).$ (2)
We assume that the point process $N=N_{\theta}$ is specified through a conditional intensity function satisfying the following assumptions.
Assumptions B.
  1. (i)
    For each $\theta\in\Theta$, $\lambda_{\theta}$ is $(\mathcal{H}_{t})_{t\in\mathbb{R}}$-predictable.
  2. (ii)
    The model $\mathscr{F}_{\Theta}$ is identifiable: $\lambda_{\theta_{1}}(0)=\lambda_{\theta_{2}}(0)$ a.s. if and only if $\theta_{1}=\theta_{2}$.
  3. (iii)
    $\lambda_{\theta}(0)\in C^{1}(\Theta)$ a.s., with for any $\theta\in\Theta$, $\partial\lambda_{\theta}^{(k)}(0)/\partial\theta_{i}\in L^{2}(\mathbb{P})$ for all $k\in[d]$ and $i\in[m]$.
  4. (iv)
    For each $\theta\in\Theta$, $\lambda_{\theta}(0)>0$ a.s.
  5. (v)
    For each $\theta\in\Theta$, there exists $\Lambda_{0}\in L^{1}(\mathbb{P})$ and a neighborhood $U=U(\theta)\subset\Theta$ of $\theta$ such that for all $\theta^{\prime}\in U$, $\|\lambda_{\theta^{\prime}}(0,\omega)\|_{\infty}\leqslant\Lambda_{0}(\omega)$.
  6. (vi)
    Let $D\lambda_{\theta}(t)=(\partial\lambda_{\theta}^{(k)}(t)/\partial\theta_{i})_{k\in[d],i\in[m]}\in\mathbb{R}^{d\times m}$ be the total derivative of $\theta\mapsto\lambda_{\theta}(t)$. Assume that
    $\alpha_{\theta}:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}D\lambda_{\theta}(t)\ \mathrm{d}t$ (3)
    exists a.s. for $\theta=\theta_{0}$.
Remark 1.
By stationarity, i.e., Assumption A(ii), Assumptions B can be stated without reference to $t$.
Remark 2.
Assumption B(vi) is satisfied if the following condition holds: there exists $\Lambda_{1}\in L^{1}(\mathbb{P})$ and a neighborhood $U=U(\theta_{0})\subset\Theta$ of $\theta_{0}$ such that for all $\theta^{\prime}\in U$, $\|D\lambda_{\theta^{\prime}}(0)\|\leqslant\Lambda_{1}$. This follows by ergodicity (Assumption A(ii)) together with the dominated convergence theorem. In this case, using stationarity of $\lambda_{\theta}$, it follows that $\alpha_{\theta}=\mathbb{E}[D\lambda_{\theta}(0)]$.
Suppose that we have some estimator $\hat{\theta}_{T}\in\Theta$ for $\theta$ based on the observation $\omega_{T}\subset[0,T]$ of $N$. We assume that this estimator satisfies a central limit theorem, as follows.
Assumption C.
There exists an $m$-variate random vector $Z$ such that under $\mathcal{P}_{\theta_{0}}$ it holds that
$\sqrt{T}(\theta_{0}-\hat{\theta}_{T})\stackrel{{\scriptstyle d}}{{\to}}Z,\quad\text{as }T\to\infty.$ (4)
Assumption C is satisfied with $Z\sim\mathcal{N}(0,I(\theta_{0})^{-1})$ (where $I(\theta)$ is defined in Assumption C.1(ii) below) when $\hat{\theta}_{T}$ is the maximum likelihood estimator and if next to Assumptions AB one grants Assumptions C.1 below. This follows from a multivariate generalization of [42], Theorem 5; see, e.g., [48], Theorem 4.
Assumptions C.1.
  1. (i)
    The log-likelihood function $L_{T}(\theta)$ at time $T$ has a unique maximum a.s. Here $L_{T}(\theta)$ is defined by
    $L_{T}(\theta)=\sum_{k=1}^{d}\left(\int_{0}^{T}\log\lambda_{\theta}^{(k)}(t)\ \mathrm{d}N^{(k)}(t)-\int_{0}^{T}\lambda_{\theta}^{(k)}(t)\ \mathrm{d}t\right).$ (5)
  2. (ii)
    For each $\theta\in\Theta$ and for all $i,j\in[m]$, $k\in[d]$, $\frac{1}{\lambda_{\theta}^{(k)}}\frac{\partial\lambda_{\theta}^{(k)}}{\partial\theta_{i}}\frac{\partial\lambda_{\theta}^{(k)}}{\partial\theta_{j}}\in L^{2}(\mathbb{P})$. Assume that for each $\theta\in\Theta$ the Fisher information matrix $I(\theta)\in\mathbb{R}^{m\times m}$ with elements
    $I_{ij}(\theta)=\sum_{k=1}^{d}\mathbb{E}\left[\frac{1}{\lambda_{\theta}^{(k)}}\frac{\partial\lambda_{\theta}^{(k)}}{\partial\theta_{i}}\frac{\partial\lambda_{\theta}^{(k)}}{\partial\theta_{j}}\right]$ (6)
    is nonsingular.
  3. (iii)
    $\lambda_{\theta}\in C^{3}(\Theta)$ with $\frac{\partial^{2}\lambda_{\theta}^{(k)}}{\partial\theta_{i}\partial\theta_{j}}\in L^{2}(\mathbb{P})$ for all $k\in[d]$, $i,j\in[m]$.
  4. (iv)
    See appendix.
Remark 3.
Assumptions A, B, and C.1 are satisfied by Poisson, Hawkes, Wold, and delayed renewal processes; see [42]. Furthermore, our setup and assumptions allow for point processes dependent on covariates, such as marked Hawkes processes.
In Section 3, we derive a limit of a certain transformation of the stochastic process $N$. We need an appropriate framework for stochastic process convergence. The transformation of the next section maps the time interval $[0,T]$ to $[0,1]$, hence it suffices to work on the latter interval. On the space $D([0,1],\mathbb{R}^{d})$ of right-continuous functions $f:[0,1]\to\mathbb{R}^{d}$ we consider multiple topologies. We consider the strong $SU$ and the weak $WU$ uniform topologies and the strong $SJ_{1}$ and weak $WJ_{1}$ Skorokhod $J_{1}$ topologies. The strong topologies are defined by treating $\mathbb{R}^{d}$ as the range space, while the weak topologies use the fact that $D([0,1],\mathbb{R}^{d})=D([0,1],\mathbb{R})^{d}$ in the sense of bijection, and consider a product topology on this space. As the name suggests, in general the strong topology is stronger than the weak topology. However, as we will prove later, the strong and product uniform topologies coincide, i.e., $SU=WU$. See [62] for more details.
3. Asymptotically distribution-free test process
In this section, we derive a distribution-free limit process from a transformation of the process $N_{\theta_{0}}$, meaning that the limit is in fact independent of the true parameter $\theta_{0}$, as well as of the model $\mathcal{F}_{\Theta}$. As indicated in the Introduction, this allows us to overcome a crucial shortcoming of goodness-of-fit tests for point processes known from the literature. Our plan of action consists of four steps.
  1. (1)
    We consider the $d$-dimensional compensated empirical process
    $\hat{\eta}^{(T)}:u\mapsto\frac{1}{\sqrt{T}}\left(N_{\theta_{0}}(uT)-\int_{0}^{uT}\lambda_{\hat{\theta}_{T}}(s)\ \mathrm{d}s\right),$ (7)
    for $u\in[0,1]$, which we decompose into
    $\hat{\eta}^{(T)}:u\mapsto\frac{1}{\sqrt{T}}\left(N_{\theta_{0}}(uT)-\int_{0}^{uT}\lambda_{\theta_{0}}(s)\ \mathrm{d}s\right)+\frac{1}{\sqrt{T}}\left(\int_{0}^{uT}\lambda_{\theta_{0}}(s)\ \mathrm{d}s-\int_{0}^{uT}\lambda_{\hat{\theta}_{T}}(s)\ \mathrm{d}s\right).$ (8)
    The first step consists of applying a martingale functional central limit theorem (martingale FCLT) to the first term of (8), in order to obtain a Gaussian limit.
  2. (2)
    We show that the second term of (8) converges to a deterministic function of $u$ multiplied by some random variable (which is, in particular, independent of $u$).
  3. (3)
    From steps 1 and 2, we obtain a limit $\hat{\eta}_{\theta_{0}}$ of (8) which is still dependent on the true parameter $\theta_{0}$, as well as on the model $\mathcal{F}_{\Theta}$. However, the particular shape of the limit allows us to apply an innovation martingale transformation $\mathscr{T}_{\theta_{0}}$, first discussed by Khmaladze in [27, 28, 29]. After transformation, $\mathscr{T}_{\theta_{0}}(\hat{\eta}_{\theta_{0}})$ is a standard Wiener process. This transformation $\mathscr{T}_{\theta_{0}}$ still depends on the true, but unknown parameter $\theta_{0}$.
  4. (4)
    Finally, we show that the difference between $\mathscr{T}_{\theta_{0}}(\hat{\eta}_{\theta_{0}})$ and its empirical counterpart $\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ is small for large $T$, hence $\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ converges to a standard Wiener process.
As we will see, due to the orthogonal structure of the martingales underlying a multivariate point process, with some work, the analysis in the multivariate case can be reduced to that of univariate case.
3.1. Step 1: FCLT for the compensated process
For our first step, we need a multivariate analog of [6], Theorem 18.3. In order to prove this, we apply the following lemma.
Lemma 1 (Cramér-Wold device for stochastic processes).
Let $(W_{n})_{n\in\mathbb{N}}\subset D([0,1],\mathbb{R}^{d})$ be a sequence of $d$-dimensional stochastic processes such that
$\text{for all }c\in\mathbb{R}^{d}:\quad c^{\top}W_{n}\stackrel{{\scriptstyle d}}{{\longrightarrow}}c^{\top}W,$ (9)
on $D([0,1],\mathbb{R})$ equipped with the Skorokhod $J_{1}$ topology, and where $W$ is a Wiener process on $\mathbb{R}^{d}$. Then it holds that $W_{n}\stackrel{{\scriptstyle d}}{{\longrightarrow}}W$ on $D([0,1],\mathbb{R}^{d})$ equipped with the strong Skorokhod $J_{1}$ topology.
Proof.
By the standard Cramér-Wold device, the condition (9) implies that the finite-dimensional distributions of $W_{n}$ converge to those of $W$.
Next, condition (9) implies that the one-dimensional marginals converge to a continuous limit, meaning that we have weak convergence in $D([0,1],\mathbb{R})$ equipped with the uniform topology. Because of this weak convergence, we have tightness (w.r.t. the uniform topology) of the one-dimensional marginals. In turn, by [62], Theorem 11.6.7, this implies tightness of the processes $W_{n}$ themselves; this tightness is w.r.t. the product topology on $D([0,1],\mathbb{R})^{d}$ obtained out of uniform topologies on $D([0,1],\mathbb{R})$. Note that here the $\|\cdot\|_{\infty}$-norm on $\mathbb{R}^{d}$ is used.
Using Prohorov’s theorem, those two conditions imply that $W_{n}\stackrel{{\scriptstyle d}}{{\longrightarrow}}W$ on $D([0,1],\mathbb{R})^{d}$ equipped with the product of uniform topologies, $WU$. We now prove that this is equal to the strong uniform topology on $D([0,1],\mathbb{R}^{d})$, $SU$.
Indeed, the product topology $WU$ is generated by the subbase consisting of the sets
$\pi_{i}^{-1}\left\{x_{i}\in D([0,1],\mathbb{R}):\|x_{i}-y_{i}\|_{\infty}<\epsilon\right\},\quad y_{i}\in D([0,1],\mathbb{R}),\epsilon>0,i\in[d],$
where $\pi_{i}$ is the projection operator onto the $i$th coordinate. Since
$\pi_{i}^{-1}\left\{x_{i}\in D([0,1],\mathbb{R}):\|x_{i}-y_{i}\|_{\infty}<\epsilon\right\}=\bigcup_{\begin{subarray}{c}z\in D([0,1],\mathbb{R}^{d})\\ z_{i}=y_{i}\end{subarray}}\left\{x\in D([0,1],\mathbb{R}^{d}):\|x-z\|_{\infty}<\epsilon\right\}$
lies in the uniform topology on $D([0,1],\mathbb{R}^{d})$, it follows that $WU\subset SU$. Next, note that SU on $D([0,1],\mathbb{R}^{d})$ is generated by the sets
$\left\{x\in D([0,1],\mathbb{R}^{d}):\|x-y\|_{\infty}<\epsilon\right\},\quad y\in D([0,1],\mathbb{R}^{d}),\epsilon>0.$
Hence, the inclusion $WU\supset SU$ follows by the equality
$\left\{x\in D([0,1],\mathbb{R}^{d}):\|x-y\|_{\infty}<\epsilon\right\}=\bigcap_{i\in[d]}\pi_{i}^{-1}\left\{x_{i}\in D([0,1],\mathbb{R}):\|x_{i}-y_{i}\|_{\infty}<\epsilon\right\}.$
To conclude, we have $W_{n}\stackrel{{\scriptstyle d}}{{\longrightarrow}}W$ on $D([0,1],\mathbb{R})^{d}$ equipped with the strong uniform topology $SU$. This implies convergence in the $SJ_{1}$ topology. ∎
Remark 4.
As is evident from the proof, Lemma 1 allows for generalizations to stochastic processes for which all the finite-dimensional distributions are Borel probability measures.
The following result is a multivariate extension of [6], Theorem 18.3.
Lemma 2 (FCLT for multivariate martingales).
Consider a stationary and ergodic two-sided $d$-variate martingale difference sequence $(\xi_{n})_{n\in\mathbb{Z}}$, where $\xi_{n}$ takes values in $\mathbb{R}^{d}$ and satisfies $\mathbb{E}[\xi_{n}|\mathcal{F}_{n-1}]=0$, with $\mathcal{F}_{n}:=\sigma(\xi_{k}:k\leqslant n)$. Write $\Xi:=(\Xi_{ij})_{i,j\in[d]}:=(\mathbb{E}[\xi^{(i)}\xi^{(j)}])_{i,j\in[d]}\in\mathbb{R}^{d\times d}$, where $\xi^{(i)}$ denotes the $i$th coordinate of the $d$-dimensional generic random variable $\xi$. The stochastic process $t\mapsto X_{t}^{n}:=\sum_{k\leqslant nt}\xi_{k}/\sqrt{n}$ converges weakly on $D([0,1],\mathbb{R}^{d})$ equipped with the strong Skorokhod $J_{1}$ topology to a $d$-variate Wiener process $W$ with covariance matrix $u\Xi$.
Proof.
By Lemma 1, it suffices to show that for arbitrary $c\in\mathbb{R}^{d}$ it holds that $c^{\top}X^{n}\stackrel{{\scriptstyle d}}{{\longrightarrow}}c^{\top}W$ on $D[0,1]$ equipped with the Skorokhod $J_{1}$ topology.
To this end, note that $(c^{\top}\xi_{n})_{n\in\mathbb{Z}}$ is a univariate martingale difference sequence w.r.t. $(\mathcal{F}_{n})_{n\in\mathbb{Z}}$, with second moment
$\mathbb{E}\left[\left(c^{\top}\xi_{k}\right)^{2}\right]=\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}\mathbb{E}\left[\xi_{n}^{(i)}\xi_{n}^{(j)}\right]=\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}\Xi_{ij}=c^{\top}\Xi c\geqslant 0.$
The limit $c^{\top}X^{n}\stackrel{{\scriptstyle d}}{{\longrightarrow}}c^{\top}W$ in the case $\mathbb{E}\left[\left(c^{\top}\xi_{k}\right)^{2}\right]=0$ is trivial, hence we may assume from now on that $\mathbb{E}\left[\left(c^{\top}\xi_{k}\right)^{2}\right]>0$. Then by [6], Theorem 18.3, it follows that $c^{\top}X^{n}$ converges weakly on $D([0,1],\mathbb{R})$ equipped with the Skorokhod $J_{1}$ topology to a Wiener process with covariance $uc^{\top}\Xi c$, i.e., to a limit $c^{\top}W$, where $W$ is a $d$-variate Wiener process with covariance $u\Xi$. ∎
The next result establishes Step 1.
Theorem 1.
Grant Assumptions A. Let $\lambda_{\theta_{0}}$ be the conditional intensity of the point process $N_{\theta_{0}}\in\mathscr{F}_{\Theta}$. Then it holds that
$\left(\frac{1}{\sqrt{T}}\left(N_{\theta_{0}}(uT)-\int_{0}^{uT}\lambda_{\theta_{0}}(s)\ \mathrm{d}s\right)\right)_{u\in[0,1]}\stackrel{{\scriptstyle d}}{{\longrightarrow}}(W_{\theta_{0}}(u))_{u\in[0,1]},$ (10)
as $T\to\infty$, weakly on $D([0,1],\mathbb{R}^{d})$ equipped with the strong Skorokhod $J_{1}$ topology, where $W_{\theta_{0}}$ is a $d$-variate Wiener process with covariance
$u\cdot\mathrm{diag}\left(\mathbb{E}[N_{\theta_{0}}^{(1)}[0,1]],\ldots,\mathbb{E}[N_{\theta_{0}}^{(d)}[0,1]]\right).$
Proof.
The result follows by an application of Lemma 2 to $\xi_{n}=N_{\theta_{0}}[n,n+1)-\int_{n}^{n+1}\lambda_{\theta_{0}}(s)\ \mathrm{d}s$, combined with an estimate of continuous-time quantities by discrete-time quantities analogous to [63], eqn. (2.19). We use the calculation
$\mathbb{E}\left[\left(N_{\theta_{0}}^{(i)}(1)-\int_{0}^{1}\lambda_{\theta_{0}}^{(i)}(s)\ \mathrm{d}s\right)\left(N_{\theta_{0}}^{(j)}(1)-\int_{0}^{1}\lambda_{\theta_{0}}^{(j)}(s)\ \mathrm{d}s\right)\right]=\begin{cases}\mathbb{E}\left[N_{\theta_{0}}^{(i)}[0,1]\right]\quad&\text{if }i=j;\\ 0\quad&\text{if }i\neq j,\end{cases}$
which follows from [16], Proposition 14.1.VIII. ∎
3.2. Step 2: Limit theorem for difference between real and empirical compensator
For the second step of our procedure, we analyze the second term of (8).
Theorem 2.
Grant Assumptions AC. Let $\alpha_{\theta_{0}}\in\mathbb{R}^{d\times m}$ be as in Assumption B(vi) and let $Z$ be as in Assumption C. Then we have, under the law $\mathcal{P}_{\theta_{0}}$ of $N_{\theta_{0}}\in\mathscr{F}_{\Theta}$,
$\left(\frac{1}{\sqrt{T}}\int_{0}^{uT}\left(\lambda_{\theta_{0}}(s)-\lambda_{\hat{\theta}_{T}}(s)\right)\ \mathrm{d}s\right)_{u\in[0,1]}\to(u\alpha_{\theta_{0}}Z)_{u\in[0,1]},$ (11)
as $T\to\infty$, weakly on $D([0,1],\mathbb{R}^{d})$ equipped with the strong Skorokhod $J_{1}$ topology.
Proof.
Our proof is structured as follows. Assumptions C grants the CLT $\sqrt{T}(\theta_{0}-\hat{\theta}_{T})\stackrel{{\scriptstyle d}}{{\to}}Z$. This allows us to perform a Taylor expansion on the integrand of the prelimit from (11): $\lambda_{\theta_{0}}(s)-\lambda_{\hat{\theta}_{T}}(s)=D\lambda_{\theta_{0}}(s)(\theta_{0}-\hat{\theta}_{T})+o_{\mathbb{P}}(|\theta_{0}-\hat{\theta}_{T}|)$. We use this to rewrite the prelimit from (11) to
$\frac{1}{\sqrt{T}}\int_{0}^{uT}\left(\lambda_{\theta_{0}}(s)-\lambda_{\hat{\theta}_{T}}(s)\right)\ \mathrm{d}s=u\frac{1}{uT}\int_{0}^{uT}D\lambda_{\theta_{0}}(s)\ \mathrm{d}s\sqrt{T}(\theta_{0}-\hat{\theta}_{T})+o_{\mathbb{P}}(1).$ (12)
We prove the result by showing that the $i$th coordinate of the right-hand side (12) converges uniformly in probability to the $i$th coordinate of the right-hand side of (11). From this, convergence in the Skorokhod topology follows.
Denote the $k$th row of $\alpha_{\theta_{0}}$ by $\alpha_{\theta_{0}}^{(k)}$, i.e.,
$\alpha_{\theta_{0}}^{(k)}:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\frac{\partial\lambda_{\theta}^{(k)}(s)}{\partial\theta}\ \mathrm{d}s\in\mathbb{R}^{m}.$
We need to show that, for each $k\in[d]$,
$u\frac{1}{uT}\int_{0}^{uT}\frac{\partial\lambda_{\theta}^{(k)}(s)}{\partial\theta}\ \mathrm{d}s\sqrt{T}(\theta_{0}-\hat{\theta}_{T})\to u\alpha_{\theta_{0}}^{(k)}Z,$ (13)
uniformly in probability. The second factor $\sqrt{T}(\theta_{0}-\hat{\theta}_{T})$ is not dependent on $u$ and converges to $Z$, by Assumptions C, hence uniform convergence in probability $\sqrt{T}(\theta_{0}-\hat{\theta}_{T})\to Z$ is immediate. Indeed, by the Skorokhod representation theorem, there exists a version of $\sqrt{T}(\theta_{0}-\hat{\theta}_{T})$ which converges in probability to a random variable with distribution $Z$.
For the first factor
$u\frac{1}{uT}\int_{0}^{uT}\frac{\partial\lambda_{\theta}^{(k)}(s)}{\partial\theta}\ \mathrm{d}s$
note that by Assumption B(iii), $\|\partial\lambda_{\theta}^{(k)}(0)/\partial\theta\|$ has finite expectation, hence by stationarity (Assumption A(ii)), the same holds for $\|\partial\lambda_{\theta}^{(k)}(s)/\partial\theta\|$, $s\geqslant 0$. In fact, it follows that $\mathbb{E}\|\partial\lambda_{\theta}^{(k)}(s)/\partial\theta\|$ is independent of time $s$. By Fubini’s theorem, we infer that
$\mathbb{E}\left[\frac{1}{T}\int_{0}^{uT}\left\|\frac{\partial\lambda_{\theta}^{(k)}(s)}{\partial\theta}\right\|\ \mathrm{d}s\right]=\frac{1}{T}\int_{0}^{uT}\mathbb{E}\left\|\frac{\partial\lambda_{\theta}^{(k)}(s)}{\partial\theta}\right\|\ \mathrm{d}s=u\mathbb{E}\left\|\frac{\partial\lambda_{\theta}^{(k)}(0)}{\partial\theta}\right\|.$
Fix $\delta,\epsilon>0$. For $u\in[0,1]$, $T>0$, an application of Markov’s inequality yields that with probability of at least $21-\delta/$ 2 it holds for all $u^{\prime}\in[0,u]$ that
$\frac{1}{T}\int_{0}^{u^{\prime}T}\left\|\frac{\partial\lambda_{\theta}^{(k)}(s)}{\partial\theta}\right\|\ \mathrm{d}s\leqslant\frac{2}{\delta}u\mathbb{E}\left\|\frac{\partial\lambda_{\theta}^{(k)}(0)}{\partial\theta}\right\|.$
Note that this bound is independent of $T>0$. Hence, selecting
$u^{(k)}=\frac{\epsilon}{4}\left(\delta\left(\mathbb{E}\left\|\frac{\partial\lambda_{\theta}^{(k)}(0)}{\partial\theta}\right\|\right)^{-1}\wedge\frac{1}{\|\alpha_{\theta_{0}}\|}\right)$
shows that with probability of at least $21-\delta/$ 2
$\sup_{u\in[0,u^{(k)}]}\left\|\frac{1}{T}\int_{0}^{uT}\frac{\partial\lambda_{\theta}^{(k)}(s)}{\partial\theta}\ \mathrm{d}s-u\alpha_{\theta_{0}}^{(k)}\right\|\leqslant\epsilon.$
Given $u^{(k)}$, we use Assumption B(vi) to select $T^{(k)}$ such that for $T\geqslant T^{(k)}$ and $u\in[u^{(k)},1]$ it holds with probability of at least $21-\delta/$ 2 that
$\left\|u\left(\frac{1}{uT}\int_{0}^{uT}\frac{\partial\lambda_{\theta}^{(k)}(s)}{\partial\theta}\ \mathrm{d}s-\alpha_{\theta_{0}}^{(k)}\right)\right\|\leqslant\epsilon.$
We have shown that on an event of probability at least $1-\delta$, for $T\geqslant T^{(k)}$ it holds that
$\sup_{u\in[0,1]}\left\|\frac{1}{T}\int_{0}^{uT}\frac{\partial\lambda_{\theta}^{(k)}(s)}{\partial\theta}\ \mathrm{d}s-u\alpha_{\theta_{0}}^{(k)}\right\|\leqslant\epsilon.$
In other words, $\frac{1}{T}\int_{0}^{uT}\frac{\partial\lambda_{\theta}^{(k)}(s)}{\partial\theta}\ \mathrm{d}s$ converges uniformly in probability to $u\alpha_{\theta_{0}}^{(k)}$.
Next, by a simple union bound similar to the one used in the proof of [59], Theorem 2.7(vi), it follows that
$\left(\frac{1}{T}\int_{0}^{uT}\frac{\partial\lambda_{\theta}^{(k)}(s)}{\partial\theta}\ \mathrm{d}s,\sqrt{T}(\theta_{0}-\hat{\theta}_{T})\right)\to\left(u\alpha_{\theta_{0}}^{(k)},Z\right),$
uniformly in probability. By an application of the continuous mapping theorem (to the function $g(x,y)=x\cdot y$) it follows that (13) holds, uniformly over $u\in[0,1]$, in probability.
We extend to (11) by selecting $u^{*}=\min_{k\in[d]}u^{(k)}$ and $T^{*}=\max_{k\in[d]}T^{(k)}$.
Note that uniform convergence in probability amounts to uniform convergence on a set of probability $21-\delta/$ 2. Therefore, conditional on a high probability set, convergence in the Skorokhod $J_{1}$ topology holds, hence the tightness conditions for the $J_{1}$ topology hold. Conditional on a set of probability $21-\delta/$ 2, the probabilities appearing in the tightness conditions can be bounded by $\delta/2$. Therefore, they can even be bounded by $\delta$. Hence, uniform convergence in probability implies convergence in the Skorokhod $J_{1}$ topology. ∎
Remark 5.
Since the limit $u\mapsto u\alpha_{\theta_{0}}Z$ is continuous, it follows that (11) even holds on $D([0,1],\mathbb{R}^{d})$ equipped with the uniform topology.
Combining Theorem 1 with Theorem 2, we have the following limit result for the empirical process $\hat{\eta}^{(T)}$ defined in (7).
Corollary 1.
Grant Assumptions AC. Let $W_{\theta_{0}}$ be a $d$-variate Wiener process with covariance
$u\cdot\mathrm{diag}\left(\mathbb{E}[N_{\theta_{0}}^{(1)}[0,1]],\ldots,\mathbb{E}[N_{\theta_{0}}^{(d)}[0,1]]\right).$
Then it holds under the law $\mathcal{P}_{\theta_{0}}$ of $N_{\theta_{0}}\in\mathscr{F}_{\Theta}$ that
$\left(\frac{1}{\sqrt{T}}\left(N_{\theta_{0}}(uT)-\int_{0}^{uT}\lambda_{\hat{\theta}_{T}}(s)\ \mathrm{d}s\right)\right)_{u\in[0,1]}\stackrel{{\scriptstyle d}}{{\longrightarrow}}(W_{\theta_{0}}(u)+u\alpha_{\theta_{0}}Z)_{u\in[0,1]}=:(\hat{\eta}_{\theta_{0}}(u))_{u\in[0,1]},$ (14)
as $T\to\infty$, weakly on $D([0,1],\mathbb{R}^{d})$ equipped with the strong Skorokhod $J_{1}$ topology.
3.3. Step 3: Transforming $\hat{\eta}_{\theta_{0}}$ into a standard Wiener process
The limiting result (14) derived in Step 2 still depends on the true but unknown parameter, through the covariance of $W_{\theta_{0}}$ and the linear stochastic drift term $u\alpha_{\theta_{0}}Z$. Due to this dependency on the true parameter, we cannot directly use this limit to construct an asymptotically valid testing procedure for all $\theta\in\Theta$, causing the theory to become fragmented.
To address this issue, we apply a suitable innovation martingale transform to the limit process, which results in a new limit process that is independent of the true parameter $\theta_{0}$ and of the specific family $\mathcal{F}_{\Theta}$ we are testing for. Variants of innovation martingale transformations have been applied to various statistical problems in the literature over the last couple of decades; see [17, 18, 30, 31, 32, 33, 40, 41, 55].
An innovation martingale transformation has also been applied to processes involving populations of size $n$ over fixed time horizons $[0,T]$, where an asymptotically distribution-free test process has been derived in the limit as $n\to\infty$ [4, 56]. However, that large-$n$ setup differs fundamentally from the large-$T$ setup considered in this work. In the large-$n$ regime, one considers a system consisting of many interacting particles over a fixed time horizon. The limiting behavior as $n\to\infty$ is typically governed by an LLN or CLT, where randomness is averaged out due to aggregation across the population. The resulting asymptotics are then driven by scaling properties of the population-level process. By contrast, the large-$T$ regime concerns a single realization of a process evolving over an increasingly long time horizon. Here, the asymptotics are dictated by the temporal structure of the process rather than by population averaging. Growth in $T$ does not inherently reduce stochasticity but instead reveals long-term dependencies and fluctuation behavior that are absent in the large-$n$ setting. As a result, techniques that rely on concentration due to population size do not carry over directly, and different probabilistic tools are required to understand the limiting behavior.
Theorem 3.
Grant Assumptions AC. Let $\mu_{\theta_{0}}^{(k)}:=\mathbb{E}[N_{\theta_{0}}^{(k)}[0,1]]$. For $\theta\in\Theta$, let $\mathscr{T}_{\theta}:D([0,1],\mathbb{R}^{d})\to D([0,1],\mathbb{R}^{d})$ be the transformation
$\mathscr{T}_{\theta}(\eta)(B):=\left\{\frac{1}{\sqrt{\mu_{\theta}^{(k)}}}\left(\eta^{(k)}(B)-\int_{B}\frac{\eta^{(k)}(1)-\eta^{(k)}(v)}{1-v}\ \mathrm{d}v\right)\right\}_{k\in[d]},\quad B\in\mathcal{B}[0,1].$ (15)
Then it holds under the law $\mathcal{P}_{\theta_{0}}$ of $N_{\theta_{0}}\in\mathscr{F}_{\Theta}$ that $\mathscr{T}_{\theta_{0}}(\hat{\eta}_{\theta_{0}})$ is a standard $d$-dimensional Wiener process, where $\hat{\eta}_{\theta_{0}}$ is defined in (14).
Proof.
Each coordinate of the limit $\hat{\eta}_{\theta_{0}}$ found in Corollary 1 is of the general form treated in [9], Section 3. Fix a coordinate $k\in[d]$. We can specify [9], Theorem 3.1, to univariate time frames by taking functions independent of the second coordinate. We identify the objects appearing in that theorem. We take the scanning family $A_{v}=[0,v)$ on $[0,1]$. We have $R(u)=\mu_{\theta_{0}}^{(k)}u$ and
$Q(B)=\int_{B}\ \mathrm{d}Q(u)=\int_{B}\ \mathrm{d}u=\mathrm{Leb}(B)=\int_{B}q(u)\ \mathrm{d}R(u)=\mu_{\theta_{0}}^{(k)}\int_{B}q(u)\ \mathrm{d}u,$
hence $q(u)=1/\mu_{\theta_{0}}^{(k)}$. Next, we have
$I(A_{v}^{c})=\int_{A_{v}^{c}}q(u)^{2}\ \mathrm{d}R(u)=\int_{v}^{1}\frac{1}{(\mu_{\theta_{0}}^{(k)})^{2}}\mu_{\theta_{0}}\ \mathrm{d}u=\frac{1-v}{\mu_{\theta_{0}}^{(k)}}.$
Note that $Q(B\cap A_{\mathrm{d}v})=Q(B\cap\{\mathrm{d}v\})=\mathbf{1}\{v\in B\}\ \mathrm{d}v$. Then [9], Theorem 3.1, gives that
$\eta_{\theta_{0}}^{(k)}(B)-\int_{B}\frac{\eta_{\theta_{0}}^{(k)}(1)-\eta_{\theta_{0}}^{(k)}(v)}{1-v}\ \mathrm{d}v$ (16)
is a Wiener process with variance $\mu_{\theta_{0}}^{(k)}u$, implying that $\mathscr{T}_{\theta_{0}}(\hat{\eta}_{\theta_{0}})$ is a standard $d$-dimensional Wiener process.
Indeed, $\mathscr{T}_{\theta_{0}}(\hat{\eta}_{\theta_{0}})$ is a $d$-dimensional Gaussian process with zero mean, hence its probabilistic behavior is characterized by its covariance structure $C_{ij}(u):=\mathrm{Cov}\left(\mathscr{T}_{\theta_{0}}(\hat{\eta}_{\theta_{0}}))_{i}(u),(\mathscr{T}_{\theta_{0}}(\hat{\eta}_{\theta_{0}}))_{j}(u)\right)$, $i,j\in[d]$, $u\in[0,1]$. Using linearity of $\mathscr{T}_{\theta_{0}}^{(i)}$ and $\mathscr{T}_{\theta_{0}}^{(j)}$ in $\eta$ together with the definition on $\hat{\eta}_{\theta_{0}}$ given in (14), we expand
$C_{ij}(u)$ $=\mathrm{Cov}\left(\mathscr{T}_{\theta_{0}}^{(i)}(W^{(i)}_{\theta_{0}}(u)),\mathscr{T}_{\theta_{0}}^{(j)}(W^{(j)}_{\theta_{0}}(u))\right)+\mathrm{Cov}\left(\mathscr{T}_{\theta_{0}}^{(i)}(W^{(i)}_{\theta_{0}}(u)),\mathscr{T}_{\theta_{0}}^{(j)}(u\alpha_{\theta_{0}}^{(j)}Z)\right)$
${\phantom{=}}\ +\mathrm{Cov}\left(\mathscr{T}_{\theta_{0}}^{(i)}(u\alpha_{\theta_{0}}^{(i)}Z),\mathscr{T}_{\theta_{0}}^{(j)}(W^{(j)}_{\theta_{0}}(u))\right)+\mathrm{Cov}\left(\mathscr{T}_{\theta_{0}}^{(i)}(u\alpha_{\theta_{0}}^{(i)}Z),\mathscr{T}_{\theta_{0}}^{(j)}(u\alpha_{\theta_{0}}^{(j)}Z))\right).$
For $i\neq j$, the first term vanishes because of independence between $W^{(i)}_{\theta_{0}}$ and $W^{(j)}_{\theta_{0}}$, while the remaining three terms vanish because $\mathscr{T}_{\theta_{0}}^{(i)}(u\alpha_{\theta_{0}}^{(k)}Z)\equiv 0$ for each $k\in[d]$. Therefore, $C_{ij}(u)\equiv 0$.∎
3.4. Step 4: Replacing $\theta_{0}$ by $\hat{\theta}_{T}$
Although the limit process from Theorem 3 does not depend on the true parameter, the transformation $\mathscr{T}_{\theta_{0}}$ and the process $(\hat{\eta}_{\theta_{0}})$ still do. Hence, in practical applications $\mathscr{T}_{\theta_{0}}(\hat{\eta}_{\theta_{0}})$ cannot be used. Instead, the best one could do is to replace it by $\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ where $\hat{\eta}^{(T)}$ is the empirical process defined in (7). Here, we show that the difference between $\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ and $\mathscr{T}_{\theta_{0}}(\hat{\eta}_{\theta_{0}})$ vanishes in the limit as $T\to\infty$. Since $\mathscr{T}_{\theta_{0}}(\hat{\eta}_{\theta_{0}})$ is a standard $d$-dimensional Wiener process, this provides us with an asymptotically distribution-free limit process. Also, the quantity $\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ can be calculated explicitly without knowledge of $\theta_{0}$.
Theorem 4.
Grant Assumptions AC. Select some $\tau\in(0,1)$. Then under the law $\mathcal{P}_{\theta_{0}}$ of $N_{\theta_{0}}\in\mathscr{F}_{\Theta}$, $\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ converges weakly on $D([0,\tau],\mathbb{R}^{d})$ equipped with the strong Skorokhod $J_{1}$-topology to a standard Wiener process, as $T\to\infty$.
Proof.
The proof consists of two parts. First, we show that $\mu_{\hat{\theta}_{T}}\to\mu_{\theta_{0}}$ as $T\to\infty$. Second, we show that
$\tilde{\mathscr{T}}:=\mathrm{diag}\left(\sqrt{\mu_{\theta_{0}}^{(1)}},\ldots,\sqrt{\mu_{\theta_{0}}^{(d)}}\right)\mathscr{T}_{\theta}$
satisfies $\tilde{\mathscr{T}}(\hat{\eta}^{(T)})\to\tilde{\mathscr{T}}(\hat{\eta}_{\theta_{0}})$ uniformly on $[0,\tau]\ni u$ in probability as $T\to\infty$; note that $\tilde{\mathscr{T}}$ could also be defined through the formula
$\tilde{\mathscr{T}}(\eta)(B):=\left\{\eta^{(k)}(B)-\int_{B}\frac{\eta^{(k)}(1)-\eta^{(k)}(v)}{1-v}\ \mathrm{d}v\right\}_{k\in[d]},\quad B\in\mathcal{B}[0,1].$
For the first part, by stationarity (Assumption A(ii)) it holds that $\mu_{\theta_{0}}^{(k)}=\mathbb{E}[\lambda_{\theta}^{(k)}(0)]$. By Assumption C, it holds that $\hat{\theta}_{T}\to\theta_{0}$ in probability, as $T\to\infty$. By passing to a subsequence, this convergence can be taken a.s. By Assumption B(iii) and the continuous mapping theorem, $\lambda_{\hat{\theta}_{T}}^{(k)}(0)\to\lambda_{\theta_{0}}^{(k)}(0)$ a.s., as $T\to\infty$. By Assumption B(v) and dominated convergence, it follows that $\mu_{\hat{\theta}_{T}}^{(k)}\to\mu_{\theta_{0}}^{(k)}$ as $T\to\infty$.
By Assumption B(iii), it holds that $\mu_{\theta_{0}}^{(k)}>0$, hence to prove the theorem it suffices to prove the second part. To this end, we apply the Skorokhod representation theorem to the result of Corollary 1 to obtain a probability space supporting probabilistic equivalent versions of $\hat{\eta}^{(T)}$ and $\hat{\eta}_{\theta_{0}}$ satisfying
[…]
We will work on this probability space.
To prove that $\tilde{\mathscr{T}}(\hat{\eta}^{(T)})\to\tilde{\mathscr{T}}(\hat{\eta}_{\theta_{0}})$ uniformly on $[0,\tau]\ni u$ in probability, it suffices to prove for $k\in[d]$ that
$\sup_{u\in[0,\tau]}\left|\int_{0}^{u}\frac{\Delta^{(T)}_{k}(1)-\Delta^{(T)}_{k}(v)}{1-v}\ \mathrm{d}v\right|\to 0$ (18)
in probability, as $T\to\infty$, i.e., to prove ucp convergence. Here $\Delta^{(T)}:=\hat{\eta}^{(T)}-\hat{\eta}_{\theta_{0}}$ denotes the difference between the empirical process from Step 2 and its limit. Since $\frac{1}{1-v}$ is bounded on $[0,\tau]$, we simply infer the convergence from (18) from (17). ∎
4. Goodness-of-fit testing procedures
4.1. Asymptotically correct goodness-of-fit test based on the test process
Given a sample $\omega_{T}$ on $[0,T]$ of a point process $T$, we are interested in testing the parametric hypothesis
$H_{0}:N\stackrel{{\scriptstyle d}}{{=}}N_{\theta}\text{ for some }\theta\in\Theta.$ (19)
In a practical situation, the sample $\omega_{T}$ would be used to calculate an estimator $\hat{\theta}_{T}$ for $\theta$, after which one asks whether $\mathscr{F}_{\Theta}=\{N_{\theta}:\theta\in\Theta\}$ is indeed a suitable model for the data. To answer this question, we use Theorem 4 to construct a goodness-of-fit test for testing the null hypothesis $H_{0}$ given in (19). The ideas of this subsection will be implemented in Sections 67. In Section 5, we consider consistency under alternatives for this goodness-of-fit test.
In Section 3, we proved in Theorem 4 that for any $\tau\in(0,1)$, the process $\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ converges weakly on $D([0,\tau],\mathbb{R}^{d})$ to a standard Wiener process $W$. The process $\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ only depends on the sample $\omega_{T}$. Therefore, it can be calculated explicitly. Now suppose $\mathscr{K}$ is a test statistic depending continuously on $\mathbb{R}^{d}$-valued stochastic processes on $[0,\tau]$. Then $\mathscr{K}(\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)}))$ converges in distribution to $\mathscr{K}(W)$.
The limit process $W$ from Theorem 4 is a standard Wiener process on $[0,\tau]$. In particular, for any $n\in\mathbb{N}$, the random variables
$Z_{i}:=\sqrt{\frac{n}{\tau}}\left(W(i\tau/n)-W((i-1)\tau/n)\right),\quad i\in[n],$ (20)
compose an i.i.d. sample of size $n$ from the $d$-dimensional standard normal distribution.
Suppose that one replaces $Z_{i}$ by $\hat{Z}^{(T)}_{i}$ based on $\hat{W}^{(T)}:=\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ instead of $W$: writing $n\equiv n(T)$, we set
$\hat{Z}^{(T)}_{i}:=\sqrt{\frac{n}{\tau}}\left(\hat{W}^{(T)}(i\tau/n)-\hat{W}^{(T)}((i-1)\tau/n)\right),\quad i\in[n].$ (21)
By the preceeding arguments, it follows that $(\hat{Z}^{(T)}_{i})_{i\in[n]}$ converges to a sample of standard normal random variables, as $T\to\infty$ in such a way that $T/n(T)\to\infty$. Hence, replacing parametric null (19) by
[…]
where $I_{d}$ denotes the $d\times d$ identity matrix, leads to aymptotically correct tests, i.e., tests having the correct rejection rates under $H_{0}$, asymptotically, as $T\to\infty$. The hypothesis (22) can be evaluated using any normality test: well-known examples include (multivariate versions of) the Kolmogorov-Smirnov, Anderson-Darling and Cramér-von Mises tests.
In order to obtain high power under alternatives, we want to take a large sample size $n=n(T)$. However, for fixed $T$, a lower value of this hyperparameter $n$ amounts to more data being used to calculate $\hat{Z}_{i}^{(T)}$, hence better approximations to standard normals, yielding a more robust testing procedure. In other words, there is a trade-off between small and large $n$. One way to find a middle ground is by taking $n=\mathrm{ceil}(c\sqrt{T})$ for some $c>0$, where $\mathrm{ceil}(\cdot)$ denotes the ceiling function. A lower choice of $c$ prioritizes the robustness of the testing procedure, while a higher $c$ prioritizes power. For this choice of the hyperparameter $n$, given a sample $\omega_{T}$ of a point process $N$ on $[0,T]$, one may test (19) using Algorithm 1.
Algorithm 1 Asymptotically correct goodness-of-fit test based on transformed empirical process
Input: A realization $\omega_{T}$ of a point process $N$ on $[0,T]$ and hyperparameters $\tau\in(0,1)$ and $n=n(T)$; e.g., use $n=\mathrm{ceil}(c\sqrt{T})$ for some $c>0$
Output: Outcome of a goodness-of-fit test
(i) Estimate $\theta$ using an estimator $\hat{\theta}_{T}$ satisfying Assumption C; e.g., use maximum likelihood estimation
(ii) Compute the compensated empirical process $\hat{\eta}^{(T)}$:
$\hat{\eta}^{(T)}:u\mapsto\frac{1}{\sqrt{T}}\left(N(uT)-\int_{0}^{uT}\lambda_{\hat{\theta}_{T}}(s)\,\mathrm{d}s\right)$
(iii) For $\mu_{\hat{\theta}_{T}}^{(k)}:=\mathbb{E}[N_{\hat{\theta}_{T}}^{(k)}[0,1]]\approx N^{(k)}(T)/T$, compute the transformed process $\hat{W}^{(T)}:=\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ using
$\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})(u):=\left\{\frac{1}{\sqrt{\mu_{\hat{\theta}_{T}}^{(k)}}}\left((\hat{\eta}^{(T)})^{(k)}(u)-\int_{0}^{u}\frac{(\hat{\eta}^{(T)})^{(k)}(1)-(\hat{\eta}^{(T)})^{(k)}(v)}{1-v}\,\mathrm{d}v\right)\right\}_{k\in[d]}$
(iv) For $i\in[n]$, compute:
$\hat{Z}^{(T)}_{i}:=\sqrt{\frac{n}{\tau}}\left(\hat{W}^{(T)}(i\tau/n)-\hat{W}^{(T)}((i-1)\tau/n)\right)$
(v) Perform an (asymptotically) exact normality test on the sample $(\hat{Z}^{(T)}_{i})_{i\in[n]}$; e.g., use the Kolmogorov-Smirnov, Anderson-Darling or Cramér-von Mises test
In the present situation, a “naive” testing procedure ignoring estimation uncertainty can be composed by simply treating the (standardized) empirical process
$\tilde{W}^{(T)}:=\mathrm{diag}\left(\sqrt{\mu_{\theta_{0}}^{(1)}},\ldots,\sqrt{\mu_{\theta_{0}}^{(d)}}\right)^{-1}\hat{\eta}^{(T)}$ (23)
as a standard Wiener process, e.g., by testing the random variables
$\tilde{Z}^{(T)}_{i}:=\sqrt{\frac{n}{\tau}}\left(\tilde{W}^{(T)}(i\tau/n)-\tilde{W}^{(T)}((i-1)\tau/n)\right),\quad i\in[n].$ (24)
for normality. This amounts to leaving out step (iii) of Algorithm 1. In Section 6, we will compare this “naive” testing procedure empirically to our asymptotically correct testing procedure, and we show empirically that it is generally a bad idea to ignore the estimation uncertainty $\hat{\theta}_{T}\neq\theta_{0}$.
Remark 6.
We know from Corollary 1 that $\tilde{W}^{(T)}$ converges to a Wiener process plus a random linear drift term. The magnitude of the bias of $\tilde{Z}_{i}^{(T)}$ caused by this drift is then of order $1/\sqrt{n}$. Since typical test statistics based on a functional of the empirical process — such as the Kolmogorov-Smirnov, Anderson-Darling and Cramér-von Mises tests — converge at rate $\sqrt{n(T)}\asymp\sqrt{T}$, the bias of the “naive” testing procedure cannot be mitigated through a choice of the hyperparameter $n$.
4.2. Random-time-change-based goodness-of-fit testing
Typically, when goodness-of-fit tests are conducted in the context of univariate point processes, one considers transformed event times $\Lambda_{\hat{\theta}_{T}}(t_{i})$, where $\Lambda_{\theta}(t)=\int_{0}^{t}\lambda_{\theta}(t)\ \mathrm{d}t$ is the compensator of the point process, i.e., the time-integrated conditional intensity. By the random time change theorem [16], Theorem 7.4.I , under $\mathcal{P}_{\theta_{0}}$ the times $\Lambda_{\theta_{0}}(t_{i})$ are distributed according to a unit Poisson process. Hence, a goodness-of-fit test is performed by testing whether $\Lambda_{\hat{\theta}_{T}}(t_{i})-\Lambda_{\hat{\theta}_{T}}(t_{i-1})$ are standard exponential. [^1] By a multivariate version of the random time change theorem, [16], Theorem 14.6.IV, this testing procedure allows for a multivariate generalization. However, the estimation uncertainty $\hat{\theta}_{T}\neq\theta_{0}$ leading to a difference between $\Lambda_{\hat{\theta}_{T}}$ and $\Lambda_{\theta_{0}}$ is typically ignored; see [3, 11, 13, 22, 34, 35, 39, 57, 58]. Even though the goodness-of-fit test outlined in Algorithm 1 relies on different ideas than this test based on the random time change theorem, it is, of course, possible to compare the outcomes of both tests.
We now consider the effect of estimation uncertainty. By Theorem 2 with $u=t_{i}/T=\mathcal{O}(1/T)$, holding for fixed $t_{i}$, it follows that, as $T\to\infty$,
$\Lambda_{\hat{\theta}_{T}}(t_{i})-\Lambda_{\hat{\theta}_{T}}(t_{i-1})=\Lambda_{\theta_{0}}(t_{i})-\Lambda_{\theta_{0}}(t_{i-1})-\frac{1}{\sqrt{T}}(t_{i}-t_{i-1})\alpha_{\theta_{0}}Z+o_{\mathbb{P}}(1/\sqrt{T}).$ (25)
By the random time change theorem, this means that the transformed interarrival times $\Lambda_{\hat{\theta}_{T}}(t_{i})-\Lambda_{\hat{\theta}_{T}}(t_{i-1})$ approximately equal standard exponential random variables plus some bias term proportial to the interarrival time $t_{i}-t_{i-1}$. This bias term leads to an incorrect goodness-of-fit testing procedure, assessing whether the times $(\Lambda_{\hat{\theta}_{T}}(t_{i}))$ compose a realization of a unit Poisson process. In fact, when one uses a test statistic that is a functional of the empirical process — such as the Kolmogorov-Smirnov, Anderson-Darling and Cramér-von Mises tests — to test the transformed interarrival times, the convergence of the tests statistics occurs at rate $\sqrt{n(T)}\asymp\sqrt{T}$. Therefore, the deviations $\frac{1}{\sqrt{T}}(t_{i}-t_{i-1})\alpha_{\theta_{0}}Z+o_{\mathbb{P}}(1/\sqrt{T})=\mathcal{O}_{\mathbb{P}}(1/\sqrt{T})$ in (25) from standard exponentials are in general not negligible.
If a consistent estimate of $\alpha_{\theta_{0}}Z$ could be obtained from the data, one could modify (25) to develop an asymptotically distribution-free goodness-of-fit test based on the random time change theorem. However, this is not a straightforward task.
5. Consistency under $H_{1}$ of the goodness-of-fit test based on the test process
In step (iii) of Algorithm 1, we transform the compensated empirical process $\hat{\eta}^{(T)}$ to $\hat{W}^{(T)}:=\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ using the innovation martingale transform defined in (7). This transformed process converges to a process independent of the true parameter $\theta_{0}$, as well as of the model $\mathscr{F}_{\Theta}$, hence leads to asymptotically correct goodness-of-fit tests. However, we notice that due to the transformation information may be lost, see [27]. In other words, it may be possible that a test based on the transformed process $\hat{W}^{(T)}$ is incapable of detecting certain deviations from the null that other tests might be able to detect. In this subsection, we discuss the consistency under $H_{1}$ of a goodness-of-fit testing procedure outlined in Algorithm 1, where for the normality test in step (v) we consider, for example, a Kolmogorov-Smirnov, an Anderson-Darling or a Cramér-von Mises test. Those tests are consistent under alternatives, meaning that if one tests a simple null $H_{0}:X\stackrel{{\scriptstyle d}}{{=}}Y$, then for a sample of $Z\stackrel{{\scriptstyle d}}{{\neq}}Y$, the power of the test converges to $1$.
Before stating the main result, we state a lemma describing the asymptotic behavior of $\hat{W}^{(T)}:=\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$.
Lemma 3.
Grant Assumptions AB and Assumptions C.1. Suppose that we observe a realization of a stationary and ergodic point process $N$ having law $\mathcal{P}$ and intensity $\lambda\notin\mathscr{L}_{\Theta}:=\{\lambda_{\theta}:\theta\in\Theta\}$; i.e., $N$ does not belong to the null (19). Let $\theta^{*}\in\Theta$ be the “least-false estimator”, or “oracle”, maximizing the expected likelihood $\mathbb{E}_{\mathcal{P}}[L_{T}(\theta)]$ over $\theta\in\Theta$. Assume that $\int_{0}^{\infty}\|\mathbb{E}[\lambda(s)-\lambda_{\theta^{*}}(s)-\mathbb{E}[\lambda(0)-\lambda_{\theta^{*}}(0)]|\mathcal{H}_{0}]\|_{2}\ \mathrm{d}s<\infty$ and $\mathrm{Var}(\lambda(0)-\lambda_{\theta^{*}}(0))<\infty$. Select $\tau\in[0,1)$.
Then the transformed empirical process $\hat{W}^{(T)}:=\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ converges to $W_{1}+W_{2}+V_{3}$, weakly on $D([0,\tau],\mathbb{R}^{d})$ equipped with the strong Skorokhod $J_{1}$-topology, where $W_{1}$ is a standard Wiener process, where $W_{2}$ is a Wiener process with covariance
$u\mapsto u\cdot\left(2\int_{0}^{\infty}\mathbb{E}\left[\frac{g_{i}(0)-\mathbb{E}[g_{i}(0)]}{\mu_{\hat{\theta}_{T}}^{(i)}}\frac{g_{j}(t)-\mathbb{E}[g_{j}(0)]}{\mu_{\hat{\theta}_{T}}^{(j)}}\right]\ \mathrm{d}t\right)_{i,j\in[d]},$ (26)
and where $V_{3}(u)=\int_{0}^{u}W_{3}(s)$, with $W_{3}$ a Wiener process also having covariance structure given by (26).
Remark 7.
The condition $\int_{0}^{\infty}\|\mathbb{E}[\lambda(s)-\lambda_{\theta^{*}}(s)-\mathbb{E}[\lambda(0)-\lambda_{\theta^{*}}(0)]|\mathcal{H}_{0}]\|_{2}\ \mathrm{d}s<\infty$ ensures that the influence of the past (before time 00) on the future (after time $t$) diminishes rapidly enough (as $t\to\infty$), meaning that the process does not have an excessively long memory.
Proof.
The maximum likelihood estimator $\hat{\theta}_{T}$ is consistent in probability for the “least-false estimator”, or “oracle”, $\theta^{*}\in\Theta$; see [23], Section 2.2. In particular, a CLT $\sqrt{T}(\theta^{*}-\hat{\theta}_{T})\stackrel{{\scriptstyle d}}{{\to}}Z$, as $T\to\infty$, still holds; cf. Assumption C.1.
Select $\tau\in(0,1)$ as in Theorem 4 and work on $[0,\tau]\ni u$. Recall that $\hat{W}^{(T)}:=\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$, where
$\hat{\eta}^{(T)}(u)$ $=\frac{1}{\sqrt{T}}\left(N(uT)-\int_{0}^{uT}\lambda(s)\ \mathrm{d}s\right)+\frac{1}{\sqrt{T}}\left(\int_{0}^{uT}\lambda(s)\ \mathrm{d}s-\int_{0}^{uT}\lambda_{\theta^{*}}(s)\ \mathrm{d}s\right)$
${\phantom{=}}\ +\frac{1}{\sqrt{T}}\left(\int_{0}^{uT}\lambda_{\theta^{*}}(s)\ \mathrm{d}s-\int_{0}^{uT}\lambda_{\hat{\theta}_{T}}(s)\ \mathrm{d}s\right).$ (27)
By the CLT $\sqrt{T}(\theta^{*}-\hat{\theta}_{T})\stackrel{{\scriptstyle d}}{{\to}}Z$ and an argument analogous to the proof of Theorem 2, the third term on the right-hand side of (27) converges to a shift term linear in $u$, with random coefficient, which is annihilated by the innovation martingale transformation $\mathscr{T}_{\hat{\theta}_{T}}$. Next, the transformation (through $\mathscr{T}_{\hat{\theta}_{T}}$) of the first term on the right-hand side of (27) converges weakly on $D([0,\tau],\mathbb{R}^{d})$ to a standard Wiener process $W_{1}$, as $T\to\infty$.
In order to be able to describe the deviation from the null $\lambda\notin\mathscr{L}_{\Theta}:=\{\lambda_{\theta}:\theta\in\Theta\}$, we consider the contribution of the second term of (27) after transformation through $\mathscr{T}_{\hat{\theta}_{T}}$. Note that
$\mathscr{T}_{\hat{\theta}_{T}}\left(u\mapsto\frac{1}{\sqrt{T}}\left(\int_{0}^{uT}\lambda(s)\ \mathrm{d}s-\int_{0}^{uT}\lambda_{\theta^{*}}(s)\ \mathrm{d}s\right)\right)$ (28)
has, after a change of variables for the second integral, $k$th component equal to
$\frac{1}{\sqrt{T}\sqrt{\mu_{\hat{\theta}_{T}}^{(k)}}}\left(\int_{0}^{uT}g_{k}(s)\ \mathrm{d}s-\int_{0}^{uT}\frac{1}{T-v}\int_{v}^{T}g_{k}(s)\ \mathrm{d}s\ \mathrm{d}v\right)=\frac{1}{\sqrt{T}\sqrt{\mu_{\hat{\theta}_{T}}^{(k)}}}\int_{0}^{uT}\left(g_{k}(s)-\bar{g}_{k}^{T}(s)\right)\ \mathrm{d}s,$ (29)
where $g_{k}(s):=\lambda^{(k)}(s)-\lambda_{\theta^{*}}^{(k)}(s)$ and $\bar{g}_{k}^{T}(s):=(T-s)^{-1}\int_{s}^{T}g_{k}(t)\ \mathrm{d}t$. Decompose (29) into
$\frac{1}{\sqrt{T}\sqrt{\mu_{\hat{\theta}_{T}}^{(k)}}}\int_{0}^{uT}\left(g_{k}(s)-\mathbb{E}[g_{k}(0)]\right)\ \mathrm{d}s+\frac{1}{\sqrt{T}\sqrt{\mu_{\hat{\theta}_{T}}^{(k)}}}\int_{0}^{uT}\left(\mathbb{E}[g_{k}(0)]-\bar{g}_{k}^{T}(s)\right)\ \mathrm{d}s.$ (30)
Using [25], Theorem VIII.3.79, with $p=q=2$ it follows that
$\frac{1}{\sqrt{T}\sqrt{\mu_{\hat{\theta}_{T}}^{(k)}}}\int_{0}^{uT}\left(g(s)-\mathbb{E}[g(0)]\right)\ \mathrm{d}s\stackrel{{\scriptstyle d}}{{\longrightarrow}}W_{2}(u),$
as $T\to\infty$, weakly on $D([0,\tau],\mathbb{R}^{d})$, where $W_{2}$ is a $d$-variate Wiener process with covariance given by (26).
Next, by the same result it follows that
$\frac{(1-u)\sqrt{T}}{\sqrt{\mu_{\hat{\theta}_{T}}^{(k)}}}\left(\mathbb{E}[g_{k}(0)]-\bar{g}_{k}^{T}(uT)\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}W_{3}^{(k)}(1-u),$
as $T\to\infty$, weakly on $D([0,1],\mathbb{R}^{d})$, for another Wiener process $W_{3}^{(k)}$ of variance (26). Apply the change of variables $s^{\prime}=Ts$ to the second term of (30). By the continuous mapping theorem [59], Theorem 18.11, applied to $h:D([0,\tau],\mathbb{R}^{d})\to D([0,\tau],\mathbb{R}^{d})$ given by $(h(f))^{(k)}(u)=\frac{1}{1-u}\int_{0}^{u}f^{(k)}(s)\ \mathrm{d}s$ it follows that the second term of (30) converges weakly on $D([0,\tau],\mathbb{R}^{d})$ equipped with the strong Skorokhod $J_{1}$-topology to $V_{3}$. ∎
Theorem 5.
Grant the conditons of Lemma 3. Fit $N$ to the parametric hypothesis $H_{0}$ by calculating the maximum likelihood estimator $\hat{\theta}_{T}$. Suppose that $\lambda$ does not satisfy
$\lambda\stackrel{{\scriptstyle d}}{{=}}\lambda_{\theta}+c\quad\text{for some}\quad\theta\in\Theta\quad\text{and some}\quad c\in\mathbb{R}^{d},$ (31)
and test the null (19) using Algorithm 1, using a goodness-of-fit test that is consistent under alternatives at rate $\sqrt{n(T)}$ in step (v). Then this goodness-of-fit test has non-trivial power, asymptotically. Furthermore, for $W_{1}$ and $W_{2}$ as in Lemma 3, suppose that
$\text{There exist }s<t,s,t\in[0,\tau]:\quad W_{1}(t)+W_{2}(t)-W_{1}(s)-W_{2}(s)\not\sim\mathcal{N}(0,t-s).$ (32)
Then the power of the test converges to $1$, as $n(T)\to\infty$.
Proof.
Lemma 3 proves convergence of $\hat{W}^{(T)}:=\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ to $W_{1}+W_{2}+V_{3}$. Note that both $W_{2}$ and $V_{3}$ are non-zero stochastic processes if there is some $k\in[d]$ such that $\mathrm{Var}(g_{k}(0))\neq 0$, hence if $\lambda$ is not equal (in $L^{1}(\mathbb{P})$) to $\lambda_{\theta}+c$ for some $\theta\in\Theta$ and some $c\in\mathbb{R}^{d}$.
Now consider the goodness-of-fit testing procedure outlined in Algorithm 1, selecting $n(T)$ such that $T/n(T)\to\infty$ and $n(T)\to\infty$ as $T\to\infty$, where we base the test on the transformed process $\hat{W}^{(T)}:=\mathscr{T}_{\hat{\theta}_{T}}(\hat{\eta}^{(T)})$ under misspecification $\lambda\notin\mathscr{L}_{\Theta}:=\{\lambda_{\theta}:\theta\in\Theta\}$. Consider step (iv) of that testing procedure. The sample $(\hat{Z}_{i}^{(T)})_{i\in[n(T)]}$ converges in distribution to a sample of standard normal random vectors coming from $W_{1}$; plus a normally distributed perturbation coming from $W_{2}$; plus a contribution coming from $V_{3}$. In order to detect the deviation from the null hypothesis, the contribution of $W_{2}+V_{3}$ should be visible in the sample $(\hat{Z}_{i}^{(T)})_{i\in[n(T)]}$, asymptotically, as the sample size $n(T)$ diverges, as $T\to\infty$.
Firstly, consider the contribution of $V_{3}$ which is defined as the integral of a Wiener process $W_{3}$, i.e., $V_{3}(t)=\int_{0}^{t}W_{3}(1-s)\ \mathrm{d}s$. Note that, with $n=n(T)$,
$\left(\sqrt{\frac{n}{\tau}}\left(V_{3}(i\tau/n)-V_{3}((i-1)\tau/n)\right)\right)_{i\in[n]}=\left(\frac{1}{\sqrt{n\tau}}W_{3}((n-i)\tau/n)+\mathcal{O}_{\mathbb{P}}(1/n)\right)_{i\in[n]}$
composes a sample of size $n(T)$, with magnitude of order $1/\sqrt{n(T)}$, having a positive autocorrelation converging to $1$ at rate $n(T)$. Typical goodness-of-fit tests based on functionals of the test process, like the Kolmogorov-Smirnov, Anderson-Darling and Cramér-von Mises tests, are consistent at rate $\sqrt{n(T)}$. Hence, the term coming from $V_{3}$, which is of magnitude $1/\sqrt{n(T)}$, contributes non-trivially to the power of the test. However, the perturbation coming from $V_{3}$ does not guarantee the power of the test to converge to $1$.
Secondly, we consider the contribution of $W_{2}$. Note that $W_{1}$ is a standard Wiener process, while $W_{2}$ is a non-zero Wiener process. When condition (32) holds, $W_{1}+W_{2}$ does not reduce to a process having variance $u\mapsto uI_{d}$. Hence, for any goodness-of-fit test in step (v) of Algorithm 1 that is consistent under alternatives (such as the Kolmogorov-Smirnov, Anderson-Darling and Cramér-von Mises tests), the deviation of $W_{1}+W_{2}$ from a standard Wiener process under the alternative $N$ is detected with probability $1$, asymptotically, as $T\to\infty$. ∎
Remark 8.
Condition (32) implicitly constrains the covariance structure between $W_{1}$ and $W_{2}$. Since $W_{1}$ is a standard Wiener process and $W_{2}$ is a non-zero Wiener process, this condition is relatively mild. For general Wiener processes $W_{1}$ and $W_{2}$ — with $W_{1}$ standard Wiener — that are dependent in a pathological way, it is possible that $W_{1}+W_{2}$ reduces to a standard Wiener process. For example, if $W_{3}$ is a standard Wiener process independent of $W_{1}$, and $W_{2}=2\rho\left(\sqrt{1-\rho^{2}}W_{3}-\rho W_{1}\right)$ for $\rho\in(0,1)$, then $W_{1}+W_{2}$ reduces to a standard Wiener process.
A natural follow-up question is whether the conclusion of Theorem 5 continues to hold under a root-$n(T)$ alternative. Since $n(T)\asymp\sqrt{T}$, this is the same as a “$T^{1/4}$ alternative”. Specifically, consider a family of alternative models $\breve{N}_{T}$ on $[0,T]$ with conditional intensity functions of the form
$\breve{\lambda}_{T}=\lambda_{\theta}+\breve{\lambda}/T^{1/4},$ (33)
where $\theta\in\Theta$ and $\breve{\lambda}$ is the conditional intensity of some other point process on $[0,T]$.
Under such alternatives, the perturbation introduced by $V_{3}$ is of order $1/n(T)$ and is therefore asymptotically undetectable. Consequently, the focus should be on the contribution of $W_{2}$. As in Theorem 5, suppose that condition (32) holds. In that case, $W_{2}$ induces a perturbation of order $1/\sqrt{n(T)}$, which leads to nontrivial power in the goodness-of-fit test. However, such root-$n(T)$ alternatives do not guarantee that the test’s power converges to one.
Corollary 2.
Use the setup of Theorem 5, but work with root-$n(T)$ alternatives. Consider the goodness-of-fit test for the null (19) using Algorithm 1, using a normality test that is consistent under alternatives at rate $\sqrt{n(T)}$ in step (v). Then this goodness-of-fit test has non-trivial power, asymptotically.
The proof of Theorem 5 implies that if the test from Section 4.1 is not consistent under an ergodic alternative with intensity $\lambda$, then it holds for some $c\in\mathbb{R}^{d}$ that $\lambda(t)-\lambda_{\theta^{*}}(t)\to c$, as $t\to\infty$. Under a stationary alternative, this is only possible if $\lambda\stackrel{{\scriptstyle d}}{{=}}\lambda_{\theta^{*}}+c$ to begin with. By choosing a model “closed under adding constant intensities” (or: a model with a parametric intensity specification including a linear parametric term), i.e., a model such that
$\{\lambda_{\theta}+c:c\in\mathbb{R}^{d},\lambda_{\theta}(s)+c\geqslant 0\text{for all }s\geqslant 0\text{ a.s.},\theta\in\Theta\}\subset\mathscr{L}_{\Theta},$
we have consistency under any stationary deviation from $H_{0}$. Note that this implies that it is a good idea to add a linear parametric term to the conditional intensity, if one wishes to ensure consistency under $H_{1}$. If our model is not “closed under adding constant intensities”, then the only stationary deviations from the null that our test might fail to detect are point processes having an intensity that is a member of
$\{\lambda_{\theta}+c:c\in\mathbb{R}^{d}\setminus\{0\},\lambda_{\theta}(s)+c\geqslant 0\text{ for all }s\geqslant 0\text{ a.s.},\theta\in\Theta\}.$
On the other hand, even for a null model “closed under adding constant intensities”, it is still possible that our test fails to detect an ergodic, but nonstationary deviation $N$ from the null, but where $\lambda\to\lambda_{\theta}$ a.s., for some $\theta\in\Theta$. However, one might argue that any such model is in fact “close” to the null, since it converges to a null model as it converges to stationarity.
6. Simulations for parametric Hawkes null hypotheses
In this section, we investigate the behavior of the goodness-of-fit tests outlined in Section 4, by considering parametric null hypotheses consisting of linear self-exciting Hawkes processes, the parameters of which we estimate from simulated sample paths of point processes. The Hawkes process was introduced in 1971, see [20, 21], and see e.g., [2, 5, 24] for more recent advances. In particular, we consider rejection rates both under simulations of point processes that are a member of the null, and of processes that are not. We compare the asymptotically exact testing procedure outlined in Algorithm 1 to “naive” testing procedures ignoring estimation uncertainty $\hat{\theta}_{T}\neq\theta_{0}$, and we argue that our asymptotically exact testing procedure has an advantage over the test based on just the compensated empirical process and over the random-time-change-based test described in Section 4.2.
For the parameter space $\Theta=(0,10)^{3}$ satisfying Assumption A(i), consider the parametric null hypothesis
$H_{0}^{\mathrm{Exp}}:N\stackrel{{\scriptstyle d}}{{=}}N_{\theta}^{\mathrm{Exp}}\text{ for some }\theta\in\{(\mu,\alpha,\beta)\in\Theta:\alpha<\beta\},$ (34)
where $N_{\theta}^{\mathrm{Exp}}=N_{\mu,\alpha,\beta}^{\mathrm{Exp}}$ is a univariate linear exponential Hawkes process with conditional intensity $\lambda_{\mu,\alpha,\beta}^{\mathrm{ExpH}}$ specified in Table 6. Consider also the similar parametric null hypothesis
$H_{0}^{\mathrm{PL}}:N\stackrel{{\scriptstyle d}}{{=}}N_{\theta}^{\mathrm{PL}}\text{ for some }\theta\in\{(\mu,\alpha,\beta)\in\Theta:\alpha<\beta\},$ (35)
where $N_{\theta}^{\mathrm{PL}}=N_{\mu,\alpha,\beta}^{\mathrm{PL}}$ is a univariate linear power-law Hawkes process with conditional intensity $\lambda_{\mu,\alpha,\beta}^{\mathrm{PLH}}$ specified in Table 6.
For both hypotheses, we simulate realizations on $[0,T]$ for $T=5{,}000$ and $T=50{,}000$. We set $n=\mathrm{ceil}(\sqrt{T}/4)$ for both the asymptotically correct transformation-based testing procedure and the “naive” testing procedure outlined in Section 4.1.
Remark 9.
The choice of the hyperparameter $n$ is motivated in Section 4.1. It turns out that $c=1/4$ yields the right rejection rates for various choices of $N_{\mu,\alpha,\beta}^{\mathrm{Exp}}$. However, one should be careful with selecting models with $\alpha$ close to $\beta$, since, for fixed $T$, for such parameters $\hat{W}^{(T)}$ might behave more erratically than a Brownian motion. This is a property inherent to the Hawkes process, not to our testing procedure; see, e.g., [26].
Using the transformation-based and the “naive” testing procedures, we perform goodness-of-fit tests for $500$ simulated realizations of the processes listed in Table 6.
Table 1. . Conditional intensities of the simulated processes
Model Name Shape of intensity
$N^{\mathrm{ExpH}}:=N_{1/2,1,2}^{\mathrm{Exp}}$ Exponential Hawkes $\lambda_{\mu,\alpha,\beta}^{\mathrm{ExpH}}:=\mu+\sum_{t_{i}<t\text{ event times of N}}\alpha e^{-\beta(t-t_{i})}$;
$N^{\mathrm{PLH}}:=N_{1/2,1,2}^{\mathrm{PL}}$ Power-law Hawkes $\lambda_{\mu,\alpha,\beta}^{\mathrm{PLH}}:=\mu+\sum_{t_{i}<t\text{ event times of N}}\alpha(1+t-t_{i})^{-\beta}$
$N^{\mathrm{SN}}_{1,2,2}$ Shot-noise $\lambda_{\mu,\alpha,\beta}^{\mathrm{SN}}:=\sum_{t_{i}<t\ \mathrm{Poi}(\mu)\text{ times}}\alpha e^{-\beta(t-t_{i})}$
$N^{\mathrm{SN}}_{1/5,10,2}$ Shot-noise $\lambda_{\mu,\alpha,\beta}^{\mathrm{SN}}:=\sum_{t_{i}<t\ \mathrm{Poi}(\mu)\text{ times}}\alpha e^{-\beta(t-t_{i})}$
$N^{\mathrm{Periodic}}:=N^{\mathrm{Periodic}}_{5/4,1,1/5,0}$ Time-inhomogeneous Poisson $\lambda_{\mu,\alpha,\beta}^{\mathrm{Periodic}}:=\mu+\alpha\sin(\beta(t-\gamma))$
$N^{\mathrm{SC}}:=N^{\mathrm{SC}}_{1,1/2,\log(2)}$ Self-correcting $\lambda_{\mu,\alpha,\beta}^{\mathrm{SC}}:=\mu e^{\beta t}\alpha^{N(t_{-})}$
Table 2. . Using the parametric null hypothesis $H_{0}^{\mathrm{Exp}}$, $T=5{,}000$ and $n=\mathrm{ceil}(\sqrt{T}/4)$, we conduct $500$ simulations for one model contained in $H_{0}$ and five models not contained in $H_{0}$. We report the number of rejections $R_{0.01};R_{0.05};R_{0.20}$ out of $500$ tests, using significance levels $0.01;0.05;0.20$, using both the transformation-based testing procedure and the “naive” testing procedure outlined in Algorithm 1, using Kolmogorov-Smirnov (KS), Cramér-von Mises (CvM) and Anderson-Darling (AD) normality tests in step (v) thereof.
Transformation-based testing procedure “Naive” testing procedure
Test KS CvM AD KS CvM AD
$N^{\mathrm{ExpH}}$ 8; 39; 108 9; 34; 116 8; 32; 123 0; 1; 6 0; 0; 7 0; 1; 12
$N^{\mathrm{PLH}}$ 7; 35; 109 5; 32; 97 9; 35; 113 0; 0; 7 0; 0; 3 0; 1; 14
$N^{\mathrm{SN}}_{1,2,2}$ 4; 23; 114 6; 27; 116 5; 26; 99 0; 0; 9 0; 0; 4 0; 0; 5
$N^{\mathrm{SN}}_{1/5,10,2}$ 7; 44; 237 0; 33; 256 0; 27; 262 0; 5; 96 0; 7; 154 0; 10; 170
$N^{\mathrm{Periodic}}$ 43; 242; 480 16; 322; 497 12; 341; 499 3; 170; 454 7; 274; 493 2; 298; 499
$N^{\mathrm{SC}}$ 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500
Table 3. . We report the number of rejections $R_{0.01};R_{0.05};R_{0.20}$ out of $500$, using significance levels $0.01;0.05;0.20$ in an experiment analogous to the one conducted for Table 6, but now with parametric null hypothesis $H_{0}^{\mathrm{Exp}}$ and $T=50{,}000$.
Transformation-based testing procedure “Naive” testing procedure
Test KS CvM AD KS CvM AD
$N^{\mathrm{ExpH}}$ 6; 25; 110 5; 20; 107 4; 25; 108 0; 0; 9 0; 0; 5 0; 0; 6
$N^{\mathrm{PLH}}$ 9; 31; 101 8; 30; 113 9; 37; 144 0; 0; 11 0; 0; 9 0; 2; 28
$N^{\mathrm{SN}}_{1,2,2}$ 6; 20; 91 5; 21; 96 4; 21; 96 0; 0; 6 0; 0; 4 0; 0; 6
$N^{\mathrm{SN}}_{1/5,10,2}$ 83; 258; 464 68; 337; 488 143; 418; 499 24; 166; 441 43; 278; 478 104; 396; 494
$N^{\mathrm{Periodic}}$ 495; 500; 500 499; 500; 500 500; 500; 500 494; 500; 500 500; 500; 500 500; 500; 500
$N^{\mathrm{SC}}$ 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500
For the simulations under $H_{0}^{\mathrm{Exp}}$, we consider the “small” time horizon $T=5{,}000$ in Table 6 and the “large” time horizon $T=50{,}000$ in Table 6. The processes (1)–(6) average around $1$ event per time unit, hence those time frames correspond to samples of around $5{,}000$ and $50{,}000$ events.
First, note that for the model $N^{\mathrm{ExpH}}$ (which belongs to $H_{0}^{\mathrm{Exp}}$), the rejection rates for the transformation-based testing procedure align well with the nominal significance levels. Specifically, we observe approximately 5, 25, and 100 rejections out of 500 at significance levels of $0.01$, $0.05$, and $0.20$, respectively. This confirms the asymptotic correctness of the transformation-based procedure, as stated in Section 4.1. In contrast, the “naive” testing procedure — based on untransformed empirical processes and ignoring the estimation uncertainty ($\hat{\theta}_{T}\neq\theta_{0}$) — produces undersized tests. For $N^{\mathrm{ExpH}}$, the number of rejections at significance levels of $0.01$, $0.05$, and $0.20$ is far below the expected counts of 5, 25, and 100, respectively. This demonstrates that ignoring estimation uncertainty leads to unreliable tests. Finally, the power of the transformation-based procedure against alternative hypotheses appears to be higher than that of the “naive” procedure. However, we emphasize that this need not hold in general.
((a))
((b))
((c))
((d))
Figure 1. . We consider $500$ simulations of $N^{\mathrm{ExpH}}$ both for $T=5{,}000$ and $T=50{,}000$, and both for the transformation-based and the “naive” testing procedures outlined in Section 4.1. We test for standard normal increments using the Kolmogorov-Smirnov test. We draw Q-Q plots for the empirical distribution of the test statistics vs. the theoretical limit distribution, i.e., the Kolmogorov distribution. The dashed red line is a straight line fitted to the scatterplot, while the line $y=x$ is drawn solid blue.
Suppose that in step (v) of the testing procedure described in Algorithm 1, we apply the Kolmogorov-Smirnov normality test. We can then compare the simulated Kolmogorov-Smirnov test statistics for $N^{\mathrm{ExpH}}\in H_{0}^{\mathrm{Exp}}$ with their theoretical null distribution given by the Kolmogorov distribution, using a Q-Q plot. Under the null, the Q-Q plot should be centered around the line $y=x$. We investigate this for $H_{0}^{\mathrm{Exp}}$ at both $T=5{,}000$ and $T=50{,}000$, comparing the transformation-based and “naive” testing procedures (see Figure 6). As expected from Sections 3 and 4.1, the Q-Q plots for the transformation-based procedure align closely with the line $y=x$ (see Figures 6(a) and (c)). Moreover, the deviation from $y=x$ is smaller at $T=50{,}000$ compared to $T=5{,}000$, which aligns with the asymptotic nature of our test as $T\to\infty$. By contrast, the “naive” testing procedure yields different results. Since the standardized empirical process in (23) corresponds to a Brownian motion with a linear drift and a random coefficient, the random variables $\tilde{Z}_{i}^{(T)}$ defined in (24) are normally distributed, but not standard normal. This is reflected in the Q-Q plots for the “naive” procedure (see Figures 6(b) and (d)), where the points form an approximately straight line but deviate from $y=x$.
Now, consider the alternative hypotheses. Detecting deviations from the null $H_{0}^{\mathrm{Exp}}$ varies in difficulty. As it turns out, the processes $N^{\mathrm{PLH}}$ and $N^{\mathrm{SN}}_{1,2,2}$ are challenging to distinguish from the null, while detecting $N^{\mathrm{SN}}_{1/5,10,2}$ is more feasible, and identifying $N^{\mathrm{Periodic}}$ is relatively easy. For $N^{\mathrm{SC}}$, the deviation from the null is straightforward to detect.
These differences can be explained as follows. First, $N^{\mathrm{PLH}}$ is a Hawkes process similar to those in $H_{0}^{\mathrm{Exp}}$, but with a power-law decay for the excitation kernel instead of exponential decay, leading to long memory. Detecting such differences in kernel shape is inherently difficult. Second, while the shot-noise process $N^{\mathrm{SN}}_{1,2,2}$ exhibits clustering behavior like the Hawkes process, it lacks self-excitation. Since its shot intensity $\mu=1$ is high compared to $\beta=2$, clusters frequently overlap, making it hard to distinguish between exogenous and endogenous clustering. By contrast, $N^{\mathrm{SN}}_{1/5,10,2}$ has a lower shot intensity than $N^{\mathrm{SN}}_{1,2,2}$, resulting in more clearly separated clusters. Since cluster initiation in the shot-noise process is exogenous, we expect the null to be more easily rejected for $N^{\mathrm{SN}}_{1/5,10,2}$ than for $N^{\mathrm{SN}}_{1,2,2}$. Now, consider the easier alternatives. The periodic Poisson process $N^{\mathrm{Periodic}}$ produces peaks in intensity, leading to concentrated event groups that resemble clusters. For small $T$, detecting this may be challenging due to the limited number of observed “clusters”. However, the regularity of these clusters should reveal that the process is not an overdispersed one. Finally, since the self-correcting process $N^{\mathrm{SC}}$ models behavior opposite to that of a Hawkes process, rejecting the null under this alternative should be quite easy.
Although Theorem 5 ensures that the power of the Kolmogorov-Smirnov, Anderson-Darling, and Cramér-von Mises tests converges to 1 as $T\to\infty$ under any alternative, the previous paragraph suggests that, in practice, detecting the deviations $N^{\mathrm{PLH}}$ and $N^{\mathrm{SN}}_{1,2,2}$ requires a sample $\omega_{T}$ over a large time frame $[0,T]$. To investigate this, we run 100 computationally intensive simulations for $H_{0}^{\mathrm{Exp}}$ versus $N^{\mathrm{PLH}}$ and $N^{\mathrm{SN}}_{1,2,2}$ for large values of $T$. The rejection rates for $N^{\mathrm{PLH}}$ are well above the significance levels at $T=5\cdot 10^{5}$, while for $N^{\mathrm{SN}}_{1,2,2}$, we need $T=5\cdot 10^{6}$ to observe similar behavior. These rejection rates are summarized in Table LABEL:tablebigT.
Table 4. . We report the number of rejections $R_{0.01};R_{0.05};R_{0.20}$ out of $100$, using significance levels $0.01;0.05;0.20$ in an experiment analogous to the one conducted for Table 6, using the parametric null hypothesis $H_{0}^{\mathrm{Exp}}$ and the alternative models $N^{\mathrm{PLH}}$ and $N^{\mathrm{SN}}_{1,2,2}$ for large time frames with $T=5\cdot 10^{5}$ and $T=5\cdot 10^{6}$, respectively.
Transformation-based testing procedure “Naive” testing procedure
Test KS CvM AD KS CvM AD
$N^{\mathrm{PLH}},T=5\cdot 10^{5}$ 3; 10; 35 2; 16; 40 3; 20; 41 0; 3; 11 0; 1; 13 0; 5; 19
$N^{\mathrm{SN}}_{1,2,2},T=5\cdot 10^{6}$ 10; 17; 40 10; 18; 41 15; 23; 42 5; 10; 29 4; 5; 21 4; 9; 22
We repeat our experiments for $H_{0}^{\mathrm{PL}}\ni N^{\mathrm{PLH}}$ consisting of univariate power-law Hawkes processes. We report our findings for $T=5{,}000$ and $T=50{,}000$ in Table LABEL:table3 and Table LABEL:table4, respectively. The observed rejection rates are comparable to the ones in Tables 6 and 6 for $H_{0}^{\mathrm{Exp}}$. This confirms the results of Sections 35 once more. Furthermore, it is possible to draw Q-Q plots analogous to Figure 6, yielding comparable results.
Table 5. . We report the number of rejections $R_{0.01};R_{0.05};R_{0.20}$ out of $500$, using significance levels $0.01;0.05;0.20$ in an experiment analogous to the one conducted for Table 6, but now with parametric null hypothesis $H_{0}^{\mathrm{PL}}$ and $T=5{,}000$.
Transformation-based testing procedure “Naive” testing procedure
Test KS CvM AD KS CvM AD
$N^{\mathrm{PLH}}$ 3; 23; 100 4; 20; 97 6; 20; 100 0; 0; 9 0; 0; 3 0; 0; 10
$N^{\mathrm{ExpH}}$ 7; 22; 92 6; 18; 90 4; 14; 86 0; 0; 4 0; 0; 2 0; 0; 3
$N^{\mathrm{SN}}_{1,2,2}$ 9; 26; 96 8; 29; 93 5; 25; 92 0; 1; 3 0; 0; 5 0; 0; 6
$N^{\mathrm{SN}}_{1/5,10,2}$ 8; 47; 243 4; 43; 298 1; 41; 304 0; 5; 140 0; 11; 204 0; 14; 245
$N^{\mathrm{Periodic}}$ 15; 131; 417 10; 169; 460 3; 173; 475 0; 46; 350 1; 98; 431 0; 111; 451
$N^{\mathrm{SC}}$ 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500
Table 6. . We report the number of rejections $R_{0.01};R_{0.05};R_{0.20}$ out of $500$, using significance levels $0.01;0.05;0.20$ in an experiment analogous to the one conducted for Table 6, but now with parametric null hypothesis $H_{0}^{\mathrm{PL}}$ and $T=50{,}000$.
Transformation-based testing procedure “Naive” testing procedure
Test KS CvM AD KS CvM AD
$N^{\mathrm{PLH}}$ 4; 24; 99 2; 19; 101 2; 18; 100 0; 0; 4 0; 0; 3 0; 0; 6
$N^{\mathrm{ExpH}}$ 3; 28; 115 4; 24; 110 2; 21; 110 0; 0; 10 0; 0; 9 0; 0; 9
$N^{\mathrm{SN}}_{1,2,2}$ 3; 22; 96 3; 22; 85 2; 20; 87 0; 0; 16 0; 0; 7 0; 0; 15
$N^{\mathrm{SN}}_{1/5,10,2}$ 104; 334; 489 123; 412; 498 244; 469; 500 43; 269; 481 87; 378; 496 204; 468; 498
$N^{\mathrm{Periodic}}$ 431; 498; 500 487; 500; 500 499; 500; 500 409; 499; 500 493; 500; 500 499; 500; 500
$N^{\mathrm{SC}}$ 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500
Remark 10.
Our results hold as well for null hypotheses different than the self-exciting ones considered here. For example, for a null hypothesis consisting of periodic Poisson processes $N^{\mathrm{Periodic}}_{\mu,\alpha,\beta,\gamma}$, we run $500$ simulations under $N^{\mathrm{Periodic}}$, with $T=5{,}000$ and using the Anderson-Darling normality test. Using significance levels $0.01$, $0.05$ and $0.20$, this yields $10$, $32$ and $104$ rejections when using the transformation-based testing procedure, and 00, 00 and $5$ rejections when using the “naive” testing procedure, respectively.
Note that the Kolmogorov-Smirnov, Anderson-Darling and Cramér-von Mises tests all yield goodness-of-fit tests of the right size when applied in step (v) of the transformation-based testing procedure outlined in Algorithm 1. Since in our simulations the power under alternatives is highest for Anderson-Darling and lowest for Kolmogorov-Smirnov, we advice to use the former in applications. We will do so in the next section.
We conclude by examining the random-time-change-based test discussed in Section 4.2, which, like the “naive” testing procedure in this section, ignores estimation uncertainty. We fit the same six processes as before to both $H_{0}^{\mathrm{Exp}}$ and $H_{0}^{\mathrm{PL}}$, using time frame $[0,T]$ for $T=5{,}000$, and then perform the goodness-of-fit test outlined in Section 4.2. The results are summarized in Table LABEL:tableGoF. We observe that the random-time-change-based test, similar to the “naive” procedure, appears undersized. Additionally, the power under any alternative seems to be lower than that of the asymptotically exact test presented in Section 4.1. Thus, the random-time-change-based test shares the shortcomings of the “naive” testing procedure.
Table 7. . We report the number of rejections $R_{0.01};R_{0.05};R_{0.20}$ out of $500$, using significance levels $0.01;0.05;0.20$ in an experiment similar to the one conducted for Table 6, but now using the testing procedure described in Section 4.2, with $T=5{,}000$.
$H_{0}^{\mathrm{Exp}}$ $H_{0}^{\mathrm{PL}}$
Test KS CvM AD KS CvM AD
$N^{\mathrm{ExpH}}$ 0; 0; 3 0; 0; 0 0; 0; 1 0; 5; 91 0; 1; 78 0; 9; 121
 $N^{\mathrm{PLH}}$ 0; 4; 49 0; 0; 36 0; 3; 59 0; 0; 1 0; 0; 1 0; 0; 2
$N^{\mathrm{SN}}_{1,2,2}$ 0; 5; 71 0; 0; 31 0; 0; 23 3; 52; 255 0; 8; 195 0; 8; 164
$N^{\mathrm{SN}}_{1/5,10,2}$ 3; 66; 239 0; 13; 212 5; 51; 271 0; 0; 7 0; 0; 2 0; 1; 9
$N^{\mathrm{Periodic}}$ 44; 165; 349 26; 163; 391 43; 193; 406 500; 500; 500 500; 500; 500 500; 500; 500
$N^{\mathrm{SC}}$ 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500 500; 500; 500
7. Data analysis
In this section, we illustrate the testing procedure outlined in Algorithm 1 by applications to real-world data. In particular, we compare the goodness-of-fit tests conducted in [44] and [54] to our asymptotically exact test.
7.1. Temporal ETAS model
First, we apply the testing procedure from Algorithm 1 to earthquake data used in [44], the seminal paper that established a framework for applying Omori’s law and the Gutenberg-Richter law to earthquake data. Building on this temporal model, subsequent years saw the development of the more advanced spatiotemporal ETAS (Epidemic Type Aftershock Sequence) model, which has become a cornerstone in earthquake modeling. For further developments and applications, see, e.g., [1, 45, 50, 60].
We consider the earthquake data set used in [44], Table 1, consisting of shallow earthquakes of less than $100$ kilometers depth, of magnitude at least $6.0$ on the Richter scale, in the Off Tohoku district east of Japan. This is the polygonal region with vertices at ($42$°N,$142$°E), ($39$°N, $142$°E), ($38$°N, $141$°E), ($35$°N, $140.5$°E), ($35$°N, $144$°E), ($39$°N, $146$°E), and ($42$°N, $146$°E); see [44], Fig. $1$ for a map. This region is one of the most seismically active areas in Japan. The data set spans the period between $1885$ and $1980$, and consists of $483$ earthquakes satisfying the depth and magnitude conditions.
The earthquake dynamics are modeled by marked point processes, with time stamps $t_{i}$ corresponding to the time of the earthquake occurrences measured in days after January 1st, 1885, and marks $m_{i}$ corresponding to earthquake magnitudes measured on the Richter scale, assumed to be independent of $t_{i}$. Since the data are recorded with a daily temporal resolution, the onset time for each earthquake was randomly assigned from a uniform distribution within the corresponding day. More specifically, [44] considers marked Hawkes processes with conditional intensity functions
$\lambda(t)=\mu+\sum_{t_{i}<t}c(m_{i})g(t-t_{i}).$ (36)
Based on the Akaike information criterion, the model with $g(t)=K/(t+c)$ and $c(m)=e^{\beta(m-M)}$ is selected. Here $M=6.0$ is the cut-off magnitude. The decay kernel $g(t)$ and the mark function $c(m)$ are theoretically motivated by Omori’s law and the Gutenberg-Richter law, respectively. This gives the conditional intensity function
$\lambda_{\theta}(t)=\lambda_{\mu,K,c,\beta}(t)=\mu+\sum_{t_{i}<t}e^{\beta(m_{i}-M)}\frac{K}{t-t_{i}+c}.$ (37)
From [44], Table 3, the maximum likelihood estimators for the parameters are
[…]
Having selected this “best” model, it is noted that it is still possible that there exists a more suitable model. To assess the suitability of the model (37), [44] considers the transformed event times $(\tau_{i})_{i\in[483]}$, where $\tau_{i}=\Lambda_{\hat{\theta}}(t_{i})$, and where $\Lambda_{\theta}(t)=\int_{0}^{t}\lambda_{\theta}(s)\ \mathrm{d}s$ is the compensator of the point process. By the random time change theorem, [16], Theorem 7.4.I, it follows that the transformed times $(\Lambda_{\theta_{0}}(t_{i}))_{i\in[483]}$ using the true parameter $\theta_{0}$ form a realization of a stationary Poisson process of unit intensity. Therefore, it is noted in [44] that if $\lambda_{\hat{\theta}}$ is a good approximation of $\lambda_{\theta_{0}}$, then $(\tau_{i})_{i\in[483]}$ is expected to behave like a stationary Poisson process of unit intensity. However, from Section 4.2 we know that even if $\lambda_{\hat{\theta}}$ is a good approximation of $\lambda_{\theta_{0}}$, the difference between $(\Lambda_{\hat{\theta}}(t_{i}))_{i\in[483]}$ and $(\Lambda_{\theta_{0}}(t_{i}))_{i\in[483]}$ is not necessarily neglectable: ignoring estimation uncertainty might lead to inaccurate tests.
In [44], the null hypothesis that $(\tau_{i})_{i\in[483]}$ is a realization of a Poisson process of unit intensity is tested with the aid of a Kolmogorov-Smirnov test; see [44], Figure 9. One could use the transformed times $(\tau_{i})_{i\in[483]}$ to construct the interarrival times $(x_{i})_{i\in[482]}$, where $x_{i}:=\tau_{i+1}-\tau_{i}$ are standard exponentially distributed under the null. Following [44], we perform a Kolmogorov-Smirnov test for the null hypothesis stating that $(x_{i})_{i\in[482]}$ is a sample from a standard exponential distribution. This yields a $p$-value of .5756.5756.5756.5756. Hence, for any reasonable significance level, the null is not rejected, indicating a seemingly good fit of the model to the data.
Figure 2. . Path of the transformed empirical process under the null hypothesis that the intensity of earthquake occurences satisfies (37). The plot exhibits deviations from the expected behavior of a standard Brownian motion, including unexpected jumps, suggesting potential inadequacies in the temporal ETAS model’s fit to the data.
We examine the situation in greater detail, accounting for the estimation uncertainty in $\hat{\theta}$ by performing the asymptotically exact transformation-based test introduced in Algorithm 1. Specifically, the transformed empirical process is calculated using (15) as outlined in step (iii) of the procedure, and its path is shown in Figure 7.1. Under the null hypothesis that the true intensity satisfies (37), this figure should resemble a standard Brownian motion. Performing the transformation-based testing procedure involves selecting a hyperparameter $n$ and a normality test for step (v) of the procedure. Unlike the setup in Section 4.1, where there is approximately one event per time unit (days), here the choice $n=\mathrm{ceil}\left(\sqrt{N[0,T]}/4\right)=6$ instead of $n=\mathrm{ceil}(\sqrt{T}/4)$ is more appropriate. Using the Anderson-Darling normality test suggested at the end of Section 6, we obtain a $p$-value of $0.0568$ (and $0.2475$ for the “naive” testing procedure).
The relatively low $p$-value and the unexpected jumps in Figure 7.1 raise concerns about the adequacy of the model’s fit. While the $p$-value does not lead to rejection at a typical significance level such as $\alpha=0.05$, it suggests that the model may not fully capture the underlying dynamics of the data. This is consistent with the expectation that the temporal ETAS model might overlook certain features better accounted for in the spatiotemporal version. As we have seen, a test based on random time change, ignoring estimation uncertainty, yields a much higher $p$-value, indicating a seemingly good fit. These discrepancies highlight that ignoring estimation uncertainty could lead to inaccurate procedures with unwarranted confidence in an otherwise imperfect model.
7.2. A recursive point process model
More recently, in 2019, a variant of the Hawkes process was introduced in [54], having the property that the expected offspring size corresponding to an event is inversely related to the conditional intensity at the time that event occurred. This “recursive point process model” is based on the idea that when a disease occurs infrequently within a population, as seen in the initial phases of an outbreak, individuals with the disease are likely to transmit it at a higher rate. In contrast, as the disease becomes more widespread, the transmission rate diminishes due to the implementation of preventive measures and the growing proportion of individuals with prior exposure. The model is described as recursive because the conditional intensity at any given moment depends on the productivity of earlier points, which itself is influenced by the conditional intensity at those same points.
This recursive model is specified through its conditional intensity
$\lambda(t)=\mu+\sum_{t_{i}<t}H(\lambda(t_{i}))g(t-t_{i}),$ (38)
with $\mu>0$, where $H:(0,\infty)\to[0,\infty)$ is assumed to be a non-increasing function and where $g:[0,\infty)\to[0,\infty)$ is a density function. Note that this model is self-exciting, and can be seen as a variant of the classical Hawkes process. A parametric specification is chosen where $H$ satisfies a power law and where $g$ is an exponential density. More specifically, $H(x)=\kappa/x^{\alpha}$ and $g(u)=\beta\exp(-\beta u)$, yielding a parametric model depending on $\theta=(\mu,\kappa,\beta,\alpha)$, where all parameters are assumed to be positive.
In [54], Section 7, this model is fitted to data consisting of $67$ recorded cases of Rocky Mountain Spotted Fever (RMSF) in California from 1960 to 2011. The data consist of time stamps marking the beginning of each week during which at least one infection occurred. Time is recorded in days since January 1st, 1960. The freely accessible data were obtained from Project Tycho (www.tycho.pitt.edu), see [47]. As in [54], Section 7, since the data are recorded with a weekly temporal resolution, the onset time for each individual case was randomly assigned from a uniform distribution within the corresponding 7-day interval. In this section, we replicate the model fitting from [54], Section 7, yielding the following parameter estimates for the California data set: [^2] A close inspection of the data set obtained from Project Tycho, in comparison with the histogram presented in [54], Figure 2, reveals a minor discrepancy between the two data sets, which cannot be attributed to a different realization of the uniform distribution within the corresponding 7-day interval. It is likely due to a correction in the data, and may account for the difference between our parameter estimates and those reported by Schoenberg.
$(\hat{\mu}_{CA},\hat{\kappa}_{CA},\hat{\beta}_{CA},\hat{\alpha}_{CA})=(0.000415,0.0775,0.00562,0.473).$ (39)
Additionally, we collect data for the state of Florida over the same time span, from 1960 to 2011, consisting of $191$ recorded disease cases. Fitting the model from [54], Section 7, to the Florida data set yields the following parameter estimates:
$(\hat{\mu}_{FL},\hat{\kappa}_{FL},\hat{\beta}_{FL},\hat{\alpha}_{FL})=(0.000777,0.833,0.0110,0.0408).$ (40)
In [54], Section 7, the model fit to the California data is assessed using super-thinned residuals. This approach is based on principles closely related to those underlying the random-time-change-based test described in Section LABEL:SectionRTC. Notably, both methods rely on plugged-in estimates for the model parameters and do not account for estimation uncertainty, which can compromise the validity of the resulting inference.
In this section, rather than using super-thinned residuals, we apply a classical goodness-of-fit test based on the random time change, as outlined in Section 4.2. After fitting the model, we transform the interarrival times $t_{i+1}-t_{i}$ using the estimated compensator, yielding the transformed values $\Lambda_{\hat{\theta}}(t_{i+1})-\Lambda_{\hat{\theta}}(t_{i})$. We then test whether these transformed interarrival times follow a standard exponential distribution. The corresponding QQ-plots, shown in Figure 7.2, include the reference line $y=x$, around which the transformed interarrival times should be centered under the null hypothesis. For the California data set, a Kolmogorov–Smirnov test yields a $p$-value of $0.3184$, which aligns with the findings in [54], Figure 3. While there is a slight deviation from the reference line $y=x$ in Figure 7.2(A), it remains within the Kolmogorov–Smirnov confidence bounds, and we therefore do not reject the null hypothesis at the $0.05$ significance level. For the Florida data set, the deviation from the reference line appears more pronounced than for the California data; nonetheless, the Kolmogorov–Smirnov test still does not reject the null hypothesis, with a $p$-value of $0.4836$.
((a))
((b))
Figure 3. . For both the California (a) and the Florida (b) data sets we fit the model for the data, after which we transform arrival times and we test for standard exponentiality of the interarrival times. There appear to be slight deviations from $y=x$, especially for the tail of the Florida data set.
Next, we account for the estimation uncertainty in $\hat{\theta}$ by applying the asymptotically exact transformation-based test described in Algorithm 1. We set the hyperparameter $n=\max\left\{\mathrm{ceil}\left(\sqrt{N[0,T]}/4\right),6\right\}=6$ for both the California data ($n([0,T])=67$) and the Florida data ($n([0,T])=191$). Following the procedure, we perform an Anderson–Darling test for normality in step (v), as suggested in Section 6.
For the California sample, this test yields a $p$-value of $0.5193$ (compared to $0.8833$ for the “naive” testing procedure). These results are consistent with the conclusions of the random-time-change-based tests above, at the $\alpha=0.05$ level. In contrast, for the Florida sample, the $p$-value is $0.0185$ (compared to $0.1979$ for the “naive” testing procedure), leading us to reject the null hypothesis when accounting for estimation uncertainty. In fact, those conclusions remain to hold if instead of an Anderson-Darling test we apply Cramér-von Mises or Kolmogorov-Smirnov tests. See Table LABEL:pvaluestab.
Although for the California data the asymptotically exact testing procedure in Algorithm 1 leads to the same conclusion as the procedures that ignore estimation uncertainty, the outcome is notably different for the Florida data: Algorithm 1 leads to a rejection of the null hypothesis, whereas the naive procedures do not. As discussed in Section 6, tests that ignore estimation uncertainty are typically undersized and have limited power under alternatives. While the model appears to fit the California data reasonably well — even though the sample size $n([0,T])=67$ is quite modest — the seemingly adequate fit suggested by the naive tests for the Florida data is contradicted by the asymptotically exact procedure. In typical settings where a new model is being evaluated, neglecting estimation uncertainty can result in misleading conclusions and overconfidence in an otherwise imperfect model. The contrasting outcomes for the Florida data set highlight the importance of accounting for estimation uncertainty to avoid erroneous inference.
Table 8. . We report $p$-values for the three goodness-of-fit testing procedures outlined in Section 4, for both the California and for the Florida data set, and using Kolmogorov-Smirnov, Cramér-von Mises or Anderson-Darling tests for standard normality or exponentiality of the obtained sample. An asterisk in the superscript denotes significance at $\alpha=0.05$.
California data Florida data
Test KS CvM AD KS CvM AD
Asymptotically exact testing procedure .5371 .4599 .5193 .0479 .0116 .0185
“Naive” testing procedure .8435 .8218 .8833 .0965 .1649 .1979
 Random-time-change based test .3104 .3319 .1895 .4836 .2770 .0759
It is important to note that the recursive model defined in (38) is self-exciting, making it a plausible choice for modeling diseases that spread directly between members of a population. However, RMSF is transmitted to humans via ticks and is not contagious between humans. Although tick bites might exhibit temporal clustering, such as spikes during favorable conditions like good hiking weather, we do not expect one tick bite to trigger another. Instead, tick bites likely follow a shot-noise type of intensity, with external factors driving exposure likelihood. As such, it is not surprising that the recursive model does not perfectly fit RMSF infection data. Given the challenges outlined in Section 6 regarding separating shot-noise from self-exciting dynamics, it is not surprising that neither the random-time-change-based test nor our asymptotically exact test rejects the null hypothesis when applied to a sample of just 67 observations.
This recursive model, however, may be particularly well-suited for diseases like COVID-19, influenza, and Ebola, which are transmissible between humans. In such scenarios, a higher number of infections (hence, a higher conditional intensity) typically triggers stricter preventive measures by policymakers or individuals, thereby decreasing the expected number of secondary infections (or the so-called “$R$-number”). This dynamic is accounted for by the recursive model but is not captured by the traditional self-exciting model.
8. Concluding remarks
In this work, we introduced an asymptotically distribution-free test process for point processes, which can be used to construct asymptotically distribution-free goodness-of-fit tests. Such tests are straightforward to implement and address the limitations of tests that ignore estimation uncertainty, which are often undersized and lack power.
Several extensions and follow-up questions arise.
  • Although our primary focus has been on $d$-dimensional point processes for $d\in\mathbb{N}$, it is worth noting that our results extend naturally to spatiotemporal point processes defined on an infinite spatial domain $\mathscr{X}$, without introducing significant challenges. In this case, one would work with a conditional intensity function $\lambda_{\theta}(s,t)$ and analyze space-integrated quantities, assuming the regularity conditions required for the central limit theorem $\sqrt{T}(\theta_{0}-\hat{\theta}_{T})\stackrel{{\scriptstyle d}}{{\to}}Z$, see [49]. Additionally, if $\alpha_{\theta}:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\int_{\mathscr{X}}\frac{\partial\lambda_{\theta}(s,t)}{\partial\theta}\,\mathrm{d}s\,\mathrm{d}t$ exists almost surely for $\theta=\theta_{0}$, one can derive a version of Corollary 1. This results in a Wiener process with variance $\mathbb{E}[N_{\theta_{0}}([0,1],\mathscr{X})]$ and drift $\alpha_{\theta_{0}}Z$, which is in the suitable form for applying an innovation martingale transformation. Therefore, we can proceed as in Section 3.3 and beyond. This framework allows for the construction of an asymptotically distribution-free goodness-of-fit test for spatiotemporal point processes.
  • We have assumed that the null hypothesis corresponds to a family $\mathscr{F}_{\Theta}$ of stationary point processes. However, in practice, one may wish to include nonstationary models, such as Hawkes processes with time-varying baseline intensities. This extension is substantially more delicate than the stationary case. Adapting our methodology would require a FCLT for nonstationary, non-ergodic martingales; that is, a nonstationary analogue of Lemma 2. Moreover, the approach relies on the availability of a central limit theorem for a suitable estimator; see [12, 36] for progress in this direction.
  • In this work, we employ the increments of the test process constructed in step (iii) of Algorithm 1. The central theoretical justification for Algorithm 1 lies in the weak convergence of the test process in step (iii) to a standard Wiener process as $T\to\infty$. This convergence permits a broad class of asymptotically valid test procedures: step (iv) may be reformulated as the application of an arbitrary continuous functional to the process in step (iii), while step (v) entails evaluating the distribution of this functional under the null via its counterpart applied to a standard Wiener process. This framework facilitates the construction of a wide range of asymptotically distribution-free goodness-of-fit tests.
  • As noted in Section 4.2, if a consistent estimate of $\alpha_{\theta_{0}}Z$ could be obtained from the data, it would be possible to modify (25) to develop an asymptotically distribution-free random-time-change-based goodness-of-fit test. However, solving this problem is far from straightforward due to the dependence of $\alpha_{\theta_{0}}Z$ on the unobservable quantity $\Lambda_{\theta_{0}}(t_{i})-\Lambda_{\theta_{0}}(t_{i-1})$ in (25). Despite this challenge, this remains an intriguing direction for future research.
Appendix
For the following set of assumptions, let $\mathcal{H}_{0,t}:=\sigma(N_{\theta}(s):s\in([0,t]),$ and let $\lambda_{\theta}^{*}$ be any $(\mathcal{H}_{t})_{t\in\mathbb{R}}$-predictable function such that
$\lambda_{\theta}^{*}(t)=\lim_{\Delta t\downarrow 0}\frac{1}{\Delta t}\mathbb{P}\left(N_{\theta}[t,t+\Delta t)>0|\mathcal{H}_{0,t}\right).$
Assumptions C.1(v).
  1. (a)
    $\mathbb{E}[\sup_{\delta\in(0,1]}\delta^{-1}(N[0,\delta])^{2}]<\infty$.
  2. (b)
    For each $\theta\in\Theta$, there exists a neighborhood $U=U(\theta)\subset\Theta$ of $\theta$ such that if
    $H(t,\omega)$ $:=\max_{\begin{subarray}{c}i,j,k\in[m]\\ \ell\in[d]\end{subarray}}\sup_{\theta^{\prime}\in U}\left|\frac{\partial^{3}\lambda_{\theta^{\prime}}^{(\ell)}}{\partial\theta_{i}\partial\theta_{j}\partial\theta_{k}}\right|;$
    $G(t,\omega)$ $:=\max_{\begin{subarray}{c}i,j,k\in[m]\\ \ell\in[d]\end{subarray}}\sup_{\theta^{\prime}\in U}\left|\frac{\partial^{3}\log\lambda_{\theta^{\prime}}^{(\ell)}}{\partial\theta_{i}\partial\theta_{j}\partial\theta_{k}}\right|,$
    then $\mathbb{E}[H(0,\omega)]<\infty$ and $\mathbb{E}[\|\lambda_{\theta_{0}}(0,\omega)\|^{2}G(0,\omega)^{2}]<\infty$.
  3. (c)
    For each $\theta\in\Theta$, $i,j\in[m]$, the following tend to zero in probability as $t\to\infty$; λ_θ-λ_θ^*, λθθi-λθ*θiand2λθθiθj-2λθ*θiθj.
  4. (d)
    For each $\theta\in\Theta$, there is some $\alpha>0$ such that the following have finite, uniformly (w.r.t. $t$) bounded $(2+\alpha)$th moments λθλθ*, 1λθ*λ*θθiλ*θθjand2λ*θθiθj, i,j[m].
  5. (e)
    For each $\theta\in\Theta$, $i\in[m]$, as $T\to\infty$
    $\mathbb{E}\left[\frac{1}{\sqrt{T}}\int_{0}^{T}\left|\frac{\partial\lambda_{\theta}}{\partial\theta_{i}}-\frac{\partial\lambda_{\theta}^{*}}{\partial\theta_{i}}\right|\ \mathrm{d}t\right]\to 0;$
    $\mathbb{E}\left[\frac{1}{\sqrt{T}}\int_{0}^{T}\left|\lambda_{\theta}-\lambda_{\theta}^{*}\right|\ \left|\frac{1}{\lambda_{\theta}^{*}}\frac{\partial\lambda^{*}_{\theta}}{\partial\theta_{i}}\right|\mathrm{d}t\right]\to 0.$
  6. (f)
    For each $\theta\in\Theta$, there exists a neighborhood $U=U(\theta)\subset\Theta$ of $\theta$ such that max_i,j,k[m]ℓ∈[d]sup_θ’U—3λθ’()(t)θiθjθk-3θ’*)()(t)θiθjθkP0 as t→∞.
  7. (g)
    For each $\theta\in\Theta$, there exists a neighborhood $U=U(\theta)\subset\Theta$ of $\theta$ and some $\alpha>0$ such that max_i,j,k[m]ℓ∈[d]sup_θ’U3θ’*)()θiθjθkandmax_i,j,k[m]ℓ∈[d]sup_θ’U3logθ’*)()θiθjθk have finite, uniformly (w.r.t. $t$) bounded $(2+\alpha)$th moments.